### v1

# Displacement correlations in a two-dimensional colloidal liquid and their relationship with shear strain correlations

Abstract
Correlations of the displacement field in a two-dimensional model colloidal liquid is studied numerically and analytically. By calculating the displacement correlations and the shear strain correlations from the numerical data of particle simulations, the displacement field is shown to have nontrivial correlations, even in liquids that are only slightly glassy with the area fraction as low as 0.50.5{0.5}0.5. It is suggested analytically and demonstrated numerically that the displacement correlations are more informative than the shear correlations: the former behaves logarithmically with regard to the spatial distance at shorter scales, while the corresponding information is missing from the shear correlations. The logarithmic behavior of the displacement correlations is interpreted as manifesting a long-lived aspect of the cage effect.
I Introduction
Dense liquids are sometimes described as “solids that flow” [1, 2]. This description, not only macroscopically but also microscopically, offers an appealing point of view, complementary to the kinetic approach to liquid states. The kinetic approach [3] is an extension of the molecular theory of gases, in which displacements do not play an essential role. In the case of dilute fluids such as gases, the molecular velocities are totally different from the hydrodynamic velocity field. In such a case, it would be a tremendous mistake to expect that the time-integral of the hydrodynamic velocity field gives the displacements of the molecule, as if the so-called “fluid parcels” were really wrapped by a membrane that confines the molecules.
For liquids of extremely high density, however, it is justifiable to define the displacement field by time-integral of the hydrodynamic velocity field [4]. As long as the liquid particles (either molecular or colloidal) are basically confined in cages, the displacement of each particle can be approximated by the value of the displacement field at its position. This is a feature in common with solids. As is well known in the case of crystalline solids, spatially correlated displacements are associated with elastic stresses. It is therefore natural to expect that, also in the cases of such high-density liquids close to vitrification (often described as “glassy”), the correlations of the displacement field may provide information about elasticity.
An interesting question is whether the displacements are correlated also in just slightly glassy liquids. Before stating this question more precisely, we need to take notice of various statistical quantities that have been devised by using the displacement as a main ingredient. Some of them are based on scalar quantities such as the density of mobile particles, while others are grounded on tensorial quantities possibly related to elasticity.
In numerical studies resolving the particle displacements in both space and time, researchers have noticed that mobile particles are distributed heterogeneously [5, 6]. The length scale of this dynamical heterogeneity has been studied typically through correlation of some quantity intended to represent the density of mobile or immobile particles, such as the magnitude of the particle displacement [7], the overlap density [8], and the self part of the intermediate scattering function [9, 10].
The formulations mentioned above are focused on the mobility of particles and not on their direction of motion. The importance of the directional aspect of motion was noticed by Doliwa and Heuer [11], who demonstrated that the displacement correlations in the longitudinal direction behave quite differently from those in the transverse direction. Later, it was noticed that the transverse displacement correlations contain information about glass elasticity [4, 12, 13]. Thus we are interested in tensorial correlations of displacements in glassy liquids, which provides a key topic of the present work.
While the displacement in solids in purely elastic response to external loading has basically smooth dependence on the spatial coordinate even at mesoscopic scales, its inelastic responses can be more localized, as is exemplified by motion of dislocations in crystalline solids. In the case of glassy materials subject to external shear, inelastic deformation occurs in the form of a localized plastic event, which, in turn, induces deformation of the surrounding elastic medium [14, 15]. The response of elastic media to such a localized inelastic source is given by the Eshelby strains [16]. Features of the Eshelby strains were reproduced from data of particle displacements by calculating correlations of mesoscopic shear strain field [17, 18, 19], which has been taken as the signature of elasticity. The Eshelby strain pattern observed with this method, interestingly, persists not only in sheared glasses behaving as elastoplastic materials but also in glassy liquids, with or without shear, even on timescales longer than the structural relaxation time.
Motivated by the above considerations, here we concretize our interest in displacement-based statistical quantities by discussing two-particle displacement correlations (DC) and shear strain correlations (SC) in two-dimensional liquids. Our first question is whether and to what extent these correlations are detectable in liquids that are only slightly dense. We focus on such density that the lifetime of “bonds” between the nearest neighbors is barely non-zero and not very long. Although the liquid in such cases does not exhibits macroscopic elasticity at all, we will show numerically an appreciable presence of DC and SC at lengthscales greater than the particle diameter. The time dependence of these correlations will also be investigated. The second question concerns the possible equivalence between the DC and the SC, as both of them are calculable from the same data of particle displacements. Does the usage of the common ingredient mean that the two kinds of correlations are equivalent? We will answer this question negatively, showing that the SC can be calculated from the DC but not vice versa. In other words, the DCs are more informative than the SCs, which leads to the third question: what is the nature of the extra information contained in DC? This question will be answered, at least partially, by showing that the DCs behave logarithmically for shorter distances, with the coefficient of the logarithmic term being an increasing function of the density. We interpret the nature of this extra information as manifestation of the cage effect.
The paper is organized as follows. We will start by reviewing some background about DC and SC and giving their definitions in Sec. II. Subsequently, in Sec. III, we will develop a theoretical framework for analytical treatment of SC; the framework, which is constructed by extending a previous work of our group on DC [20], allows us to express SC in terms of DC. After specifying the two-dimensional Brownian particle system in Sec. IV as a model liquid, we will compute the DC and the SC from the numerical data in Sec. V. As a result, the DCs will be shown to behave logarithmically at lengthscales between the particle diameter and the correlation length, which makes the DC more sensitive to the mean density (measured by the area fraction) than the SC. Possible origin of the logarithmic behavior of DC is discussed in Sec. VI, along with the fact that the DC and the SC can be expressed in terms of a similarity variable with diffusive dynamical length. Finally, Sec. VII is devoted to concluding remarks.
II Background
II.1 Caged dynamics in liquids as “solids that flow”
As a background to the questions about DCs and SCs in liquids, we begin this review section with the notion of cage effects in dense liquids regarded as solids with impermanent bonds.
We have noted in Introduction that the approach from the solid-like aspect of liquids is complementary to that by extending the kinetic theory of gases toward higher densities. The kinetic approach [3] seems more naturally applicable when the correlations between collisions are smaller; this occurs if the density is so low that collision is basically a process in which a molecule encounters a new partner every time. In other words, reunion with the same molecule (as well as ring collision) is a rare event in dilute fluids.
In dense liquids, contrastively, neighboring particles stay neighbors for a long time. This occurs even in the absence of attractive interactions. Every particle tends to remain in contact with the same neighbors, often described as a cage in which the particle is trapped. Although the notion of repeated collisions is still applicable to the interaction between the caged particle and its neighbors, the effect of ceaseless interaction between the same pair of particles is more appropriately regarded as a kind of bond with a long but finite lifetime [5, 21]. In the limiting case of permanent bonds, the material would behave as a solid (either crystalline or amorphous).
The presence of “bond breakage” as a rare event introduces fluidity. The lifetime of bonds, in the absence of external shear, is on the order of the structural relaxation time $\tau_{\alpha}$, conventionally defined as the timescale of the intermediate scattering function at the scale of the particle diameter [22].
Thus the microscopic dynamics of dense liquids for timescales shorter than $\tau_{\alpha}$ are solid-like, in the sense that they are dominated by such “bonds” with a long lifetime. The total effect of such a bond structure could be conceived as confinement by nested cages, and therefore often referred to as the cage effect. For timescales longer than $\tau_{\alpha}$, the “bonds” are broken and therefore the dynamics cannot be purely elastic, but we note the possibility that the collapse of the whole cage structure may take much longer time than $\tau_{\alpha}$. The dynamics are then expected to be viscoelastic, possibly depending on lengthscales in some intriguing manner.
II.2 Displacement in liquids
Let us continue reviewing statistical physics of dense liquids in general, which can be molecular or colloidal for the present.
The cage effect suggests the possibility that, in dense liquids, hydrodynamic description may be extended to shorter lengthscales close to the interparticle distance [3, 23]. Hydrodynamics means finding a closed set of dynamical equations for slow variables. The long lifetime of the “bonds” can provide grounds for slow variables in dense liquids, such as the temporally coarse-grained velocity of a particle. We emphasize that the situation is different from that of low-density fluids, in which the hydrodynamic velocity field is introduced by spatial coarse-graining of momentum as a conserved quantity. In deriving the Navier–Stokes equation for Newtonian fluids, any confusion between the velocity of a fluid element and the molecular velocity should be severely criticized. In dense liquids, however, the cage effect makes it possible—probably as a crude approximation—to regard the velocity of a particle as the velocity of a “fluid element” comprising the particle, with the understanding that fluctuations at the collisional frequency should be filtered out mostly by time-averaging.
A simple way to remove the fast fluctuations in the velocity is to focus on the displacement of the particle [24]. Let us denote the position vector of the $i$-th particle with $\mathbf{r}_{i}=\mathbf{r}_{i}(t)$; and its displacement, for the time interval from $s$ to $t$, with $\mathbf{R}_{i}=\mathbf{R}_{i}(t,s)$. The displacement of the $i$-th particle is then given by integrating the velocity $\dot{\mathbf{r}}_{i}$ over the time interval, as
$\mathbf{R}_{i}(t,s)=\mathbf{r}_{i}(t)-\mathbf{r}_{i}(s)=\int_{s}^{t}{\dot{\mathbf{r}}_{i}(t^{\prime})}\,\mathrm{d}{t^{\prime}}.$ (1)
The most fundamental one-particle quantity based on $\mathbf{R}_{i}(t,s)$ is the mean square displacement (MSD), which we write symbolically as $\left\langle{\mathbf{R}^{2}}\right\rangle$ [25]. On the assumption that the system is statistically steady, the MSD for ${t-s}\gg\tau_{\alpha}$ grows asymptotically in proportion to $t-s$. This asymptotic behavior is understood in terms of “steps” at $t_{m}=s+m\tau_{\alpha}$ ($m=0,1,\ldots$), defined by $\mathbf{R}_{i}(t_{m+1},t_{m})$, whose accumulation gives
$\mathbf{R}_{i}(t_{M},t_{0})=\sum_{m=0}^{M-1}\mathbf{R}_{i}(t_{m+1},t_{m});$ (2)
if separate steps are uncorrelated, the MSD for the time interval from $t_{0}$ to $t_{M}$ (with $M$ being a positive integer) is proportional to $M=(t_{M}-t_{0})/\tau_{\alpha}$.
Statistical quantities based on displacements of two or more particles are indicative of collective motions related to various aspects of the caged dynamics [6, 8, 26]. In the presence of such collective motions, in which the displacements of distinct particles are spatially correlated, it seems justifiable to define a displacement field in some way. Thus the hydrodynamical description is extended to shorter scales without requiring momentum conservation. The existence of the displacement field allows us to study the “solidity” of the liquid by comparing the behavior of its displacement field with that of viscoelastic continuum models.
II.3 Displacement correlations
From among various statistical quantities based on displacements in liquids, here we focus on one of its simplest form involving two particles, referring to it as the (two-particle) displacement correlation (DC) tensor. In terms of the particle displacement, given in Eq. (1), the DC tensor is defined as follows [27, 20]: Denoting the Cartesian components of the displacement vector with
$\mathbf{R}_{i}=\begin{bmatrix}R_{ix}\\ R_{iy}\end{bmatrix}=\begin{bmatrix}R_{ix}(t,s)\\ R_{iy}(t,s)\end{bmatrix}$ (3)
in the 2D setup (and omitting the time arguments when obvious), we consider the tensorial product of displacements of two particles ($i$ and $j$),
$\mathbf{R}_{i}\otimes\mathbf{R}_{j}=\begin{bmatrix}{R_{ix}}{R_{jx}}&{R_{ix}}{R_{jy}}\\ {R_{iy}}{R_{jx}}&{R_{iy}}{R_{jy}}\end{bmatrix}.$ (4)
The DC tensor is then defined by averaging $\mathbf{R}_{i}\otimes\mathbf{R}_{j}$ over all pairs $(i,j)$ such that their relative position vector, at the “initial” time $s$, equals a given vector $\tilde{\mathbf{d}}$. On the assumption that the system is statistically homogeneous, hereafter we denote the DC tensor with a capital $\chi$ as $\mathsf{X}(\tilde{\mathbf{d}},t,s)$, making it clear that the independent variables are $\tilde{\mathbf{d}}$, $t$ and $s$; then we write its definition symbolically as
$\mathsf{X}(\tilde{\mathbf{d}},t,s)=\left\langle{\mathbf{R}\otimes\mathbf{R}}\right\rangle_{\tilde{\mathbf{d}}}=\begin{bmatrix}\left\langle{{R_{x}}{R_{x}}}\right\rangle_{\tilde{\mathbf{d}}}&\left\langle{{R_{x}}{R_{y}}}\right\rangle_{\tilde{\mathbf{d}}}\\ \left\langle{{R_{y}}{R_{x}}}\right\rangle_{\tilde{\mathbf{d}}}&\left\langle{{R_{y}}{R_{y}}}\right\rangle_{\tilde{\mathbf{d}}}\end{bmatrix},$ (5)
where $\left\langle{\quad}\right\rangle_{\tilde{\mathbf{d}}}$ denotes conditional average over the pairs $(i,j)$ satisfying $\mathbf{r}_{j}(s)-\mathbf{r}_{i}(s)=\tilde{\mathbf{d}}$; see Eq. (84) in Appendix A. For further details, see also Subsec. III-A of Ref. [20].
On the assumption of statistical steadiness, the DCs depend on the time interval only through $t_{\Delta}=t-s$. Besides, isotropy and reflectional symmetry are also assumed, so that the DC tensor must be decomposable into the longitudinal and the transverse correlations, denoted with $X_{\parallel}$ and $X_{\perp}$, respectively. We write this decomposition as
[…]
where $\tilde{{d}}=\left|{\smash{\tilde{\mathbf{d}}}}\right|$. Numerical procedure for calculating these correlations from particle simulation data is explained in Appendix A.
(a)Refer to caption(b)Refer to caption
Figure 1: Typical behavior of DCs. The numerical conditions are $\phi_{\text{area}}=0.5$ and $t_{\Delta}=20\tau_{0}$ (the same as in Fig. 5 shown later); see the main text for details. (a) Longitudinal and transverse DCs, plotted against initial distance $\tilde{{d}}$. The particle diameter $\sigma$ is taken as a reference length scale. (b) The numerical value of $\mathsf{X}(\tilde{\mathbf{d}},t_{\Delta})\cdot\mathbf{e}_{1}$, plotted as a vector field on the $\tilde{\mathbf{d}}$-plane (in the fashion of Fig. 1 in Ref. [20], with $\mathbf{e}_{1}$ denoting the ${x}$-directional unit vector). The axes are normalized with the diffusive length $2\sqrt{D^{\text{c}}{t}}$ appearing in Eq. (50).
Typical behavior of the DCs as functions of $\tilde{\mathbf{d}}$, with $t_{\Delta}$ fixed, is depicted in Fig. 1. Details of the particle model used here [20] will be specified later in Sec. IV. It is seen in Fig. 1(a) that the two components of DC behave quite differently: $X_{\parallel}$ is positive everywhere, while $X_{\perp}$ becomes negative for large $\tilde{{d}}$. This difference was noted in a pioneering work by Doliwa and Heuer [11], many years before the researchers started to notice that the transverse DC may contain information about shear modulus of glassy systems [13, 12]. The directional behavior of DCs can be illustrated pictorially as a pair of vortices [20, 11], as is exemplified in Fig. 1(b).
Turning our attention to the $t_{\Delta}$-dependence of DCs, we find a rather curious fact: in comparison to $\tau_{\alpha}$ determined as the timescale of $\left\langle{e^{\mathrm{i}\mathbf{k}\cdot\mathbf{R}}}\right\rangle$, DCs are quite long-lived. In the case of the “vortices” reported by Doliwa and Heuer [11], in a 2D system of Brownian particles (disks) with the area fraction $\phi_{\text{area}}=0.77$, the time interval chosen for their Fig. 8 was as long as $10\tau_{\alpha}$. The same behavior of $X_{\parallel}$ and $X_{\perp}$ is observed even for colloidal liquids which are only slightly glassy. In the case of $\phi_{\text{area}}=0.5$ studied in Ref. [20], the area fraction is so small that $\tau_{\alpha}$ is no longer than the microscopic time scale $\tau_{0}=\sigma^{2}/D$ (composed of the disk diameter $\sigma$ and the bare diffusivity $D$; see Sec. IV), and yet the displacements are correlated for much longer timescales, as was shown in Fig. 1 of Ref. [20] ($t_{\Delta}=0.8\,\tau_{0}$), which is now corroborated in Fig. 1 of the present article ($t_{\Delta}=20\tau_{0}$).
To explain the non-vanishing DC for $t_{\Delta}\gg\tau_{\alpha}$, Doliwa and Heuer [11] argued that interparticle correlations from shorter times can still contribute to DC for $t_{\Delta}$. If the two-particle displacements are correlated only within $\tau_{\alpha}$, we can show that the contributions to DC are accumulated in proportion to $t_{\Delta}$, by following the same line of argument as that for MSD with Eq. (2). We will demonstrate in Sec. V, however, that the actual $t_{\Delta}$-dependence of DC differs from the prediction of this simple scenario. The numerical result suggests that some aspects of the cage structure, probed by DC, are more long-lived than $\tau_{\alpha}$.
II.4 Shear strain correlations
The directionality and longevity of the DCs are reminiscent of the behavior of mesoscopic shear strains [17, 18, 19], whose correlations has been recognized as the Eshelby strain pattern [16, 14] depicted as a four-petaled flower-like figure. We refer to this kind correlations as shear strain correlations or, more concisely, shear correlation (SC).
The four-petaled SC pattern is often interpreted as a sign of elasticity, showing the response of an elastic media to localized plastic events. Although the observations of the SC pattern in sheared glasses [17, 18] seems to be consistent with this interpretation, it is curious that non-vanishing SC is reported also in quiescent glassy liquids, being visible even for timescales longer than $\tau_{\alpha}$ [19].
Before studying the behavior of SC for $t_{\Delta}\gg\tau_{\alpha}$, here we review how to calculate the SC from the particle simulation data. A crucial assumption for definition of the SC is the existence of the displacement field. Considering that there can be some choice of independent variables in continuum mechanics, here we choose $\mathbf{r}(s)$, the position of the “fluid element” at the time $s$, as the spatial independent variable. Following basically the procedure given by Illing et al. [19] in their supplementary material, we define the displacement field $\mathbf{R}=\mathbf{R}(\mathbf{r}(s),t,s)$ as
$\mathbf{R}(\mathbf{r}(s),t,s)=\left.{\frac{1}{\rho(\mathbf{x},s)}\sum_{i}\mathbf{R}_{i}(t,s)\bar{\delta}_{\varepsilon}(\mathbf{r}_{i}(s)-\mathbf{x})}\right|_{\mathbf{x}=\mathbf{r}(s)},$ (7)
where $\bar{\delta}_{\varepsilon}(~{}\cdot~{})$ is a slightly blurred delta function [28], and $\rho$ is the density field defined as
$\rho(\mathbf{x},t)=\sum_{i}\bar{\delta}_{\varepsilon}(\mathbf{r}_{i}(t)-\mathbf{x}).$ (8)
Taking the statistical steadiness into account, we can write $\mathbf{R}(\mathbf{r}(s),t,s)=\mathbf{R}(\mathbf{r}(s),t_{\Delta})$. Once $\mathbf{R}$ is defined, the relative deformation for the time interval from $s$ to $t$ is given by the mapping
$\mathbf{r}(s)=\begin{bmatrix}x(s)\\ y(s)\end{bmatrix}\mapsto\mathbf{r}(t)=\begin{bmatrix}x(t)\\ y(t)\end{bmatrix}=\mathbf{r}(s)+\mathbf{R}(\mathbf{r}(s),t_{\Delta}).$ (9)
The deformation gradient relative to the configuration at the time $s$, which we denote with $\mathsf{F}$, is then given by
$\mathsf{F}=\frac{\partial\mathbf{r}(t)}{\partial\mathbf{r}(s)}=\begin{bmatrix}\partial{x(t)}/\partial{x(s)}&\partial{x(t)}/\partial{y(s)}\\ \partial{y(t)}/\partial{x(s)}&\partial{y(t)}/\partial{y(s)}\end{bmatrix}.$ (10)
The symmetric part of $\mathsf{F}$ gives the strain tensor. In particular, the sum of the off-diagonal components, which we denote with
$\gamma=\gamma(\mathbf{r}(s),t,s)=F_{12}+F_{21}=\frac{\partial x(t)}{\partial y(s)}+\frac{\partial y(t)}{\partial x(s)},$ (11)
would correspond to the shear strain if the system would be driven along the $x$-axis so that $\left\langle{\mathbf{R}(\mathbf{r}(s),t_{\Delta})}\right\rangle\propto(y(s)t_{\Delta},0)$. In the absence of external driving, still we refer to $\gamma$ in Eq. (11) as the shear strain field. The SC is defined as correlation of $\gamma$ at two positions separated by $\tilde{\mathbf{d}}$ (say, at $\mathbf{x}^{\prime}$ and $\mathbf{x}^{\prime}+\tilde{\mathbf{d}}$) [25]:
$\chi_{\gamma}=\left\langle{\gamma\gamma}\right\rangle_{\tilde{\mathbf{d}}}=\left\langle{\gamma({\mathbf{x}^{\prime}},t,s)\gamma({\mathbf{x}^{\prime}}+\smash{\tilde{\mathbf{d}}},t,s)}\right\rangle.$ (12)
Due to the spatial and temporal translational symmetry of the system, $\chi_{\gamma}$ depends on the space–time interval $(\tilde{\mathbf{d}},t_{\Delta})$ alone, and not on $\mathbf{x}^{\prime}$ and $s$ directly.
Figure 2: Shear correlation $\chi_{\gamma}$ given by Eq. (63) with ${\mu_{\text{r}}}=0.25$. On the understanding that $\vartheta\propto\xi_{*}\propto\left|{\smash{\tilde{\mathbf{d}}}}\right|$, the value of $\chi_{\gamma}/(A\tilde{\mathbf{d}}^{2})$ is plotted as a color map on the $\tilde{\mathbf{d}}$-plane, being normalized with the factor $A$ to fit within the range from $-1$ to $+1$. The axes are normalized with the diffusive length $2\sqrt{D^{\text{c}}{t}}$ appearing in Eq. (50).
In order to answer the questions about the correlations raised in Introduction, we will make comparative studies of DC and SC in the following sections. In particular, a formula interconnecting the DC and the SC will be developed in Sec. III. As a usage demonstration of the formula, we can obtain an approximate analytical expression for SC, later given as Eq. (63), from that of DC given in Ref. [20]. The result is shown in Fig. 2 as a color map on the $\tilde{\mathbf{d}}$-plane, in which the four-petal flower-like pattern is evident.
III Relations involving shear strain correlations
As we reviewed in the previous section, the DC and the SC share a common ingredient, in the sense that both are calculated from the displacements of the particles. The two kinds of correlations have something in common, for example, in that they reflect the vectorial character of the displacement, and also in that they can persist longer than $\tau_{\alpha}$.
These observations have led us to ask whether the two correlations are equivalent to each other. Although the displacement is used as a common ingredient for the DC and the SC, we have a reason to doubt their equivalence: there is an evident characteristic length in the transverse DC in Fig. 1, marked by the point at which $X_{\perp}$ changes the sign or by the minimum of $X_{\perp}$, while such a length scale is not so apparent in SC.
In order to answer this question, let us develop a theoretical framework to derive an analytical relation between the DC and the SC.
III.1 Label variable and deformation gradient tensor field
In order to relate the DC with SC analytically, we need to formulate the two kinds of correlations on the same ground, taking the analytical amenability into account. For this aim, we adopt the label variable formulation developed in our previous works on the DC tensor [27, 20].
The basic idea is to introduce the label variable $\bm{\xi}=(\xi_{1},\xi_{2})$ convected by the velocity field $\mathbf{u}$, which means that $\xi_{\alpha}=\xi_{\alpha}(\mathbf{x},t)$ ($\alpha\in\{1,2\}$) satisfies
$\left(\partial_{t}+\mathbf{u}\cdot\frac{\partial}{\partial\mathbf{x}}\right)\xi_{\alpha}(\mathbf{x},t)=0$ (13)
where
$\mathbf{u}=\mathbf{u}(\mathbf{x},t)=\frac{\mathbf{Q}(\mathbf{x},t)}{\rho(\mathbf{x},t)},\quad\mathbf{Q}(\mathbf{x},t)=\sum_{i}\dot{\mathbf{r}}_{i}(t)\bar{\delta}_{\varepsilon}(\mathbf{r}_{i}(t)-\mathbf{x}),$ (14)
and $\rho(\mathbf{x},t)$ is the density field given by Eq. (8). From among infinitely many solutions to Eq. (13) corresponding to different initial data, we choose a solution that satisfies $\partial(\xi_{1},\xi_{2})/\partial(x_{1},x_{2})=\rho(\mathbf{x},t)$ with the aid of the continuity equation [27], thus constructing $\bm{\xi}=\bm{\xi}(\mathbf{x},t)$ subject to Eq. (13) at each instant and normalized so that
$\left|{\bm{\xi}(\mathbf{r}_{i}(t),t)-\bm{\xi}(\mathbf{r}_{j}(t),t)}\right|\simeq\frac{\left|{\mathbf{r}_{i}(t)-\mathbf{r}_{j}(t)}\right|}{\ell_{0}};$
here $\ell_{0}=1/\sqrt{\rho_{0}}$ is the typical interparticle distance, defined in terms of the mean density $\rho_{0}$. Then, by taking $\bm{\xi}$ as the independent variable, we transfer to a $t$-dependent curvilinear coordinate system sticking to the particles (known by the name of the convective coordinate system), in the form of a mapping
$\bm{\xi}\mapsto\mathbf{r}=\mathbf{r}(\bm{\xi},t)=\begin{bmatrix}x\\ y\end{bmatrix}$ (15)
such that $\mathbf{r}(\bm{\xi}(\mathbf{x},t),t)=\mathbf{x}$.
While the partial derivative of $\mathbf{r}=\mathbf{r}(\bm{\xi},t)$ with regard to $t$ gives the velocity $\mathbf{u}$, its partial derivative with regard to $\bm{\xi}$ yields what is called the deformation gradient tensor [29] or the displacement gradient tensor [30]. It is convenient to rearrange the components of this tensor field, $\partial\mathbf{r}/\partial\bm{\xi}$, into a form corresponding to the Helmholtz decomposition of the displacement field [4, 20], which consists of the dilatational component ${\Psi}_{\mathrm{d}}$ and the rotational component ${\Psi}_{\mathrm{r}}$. We write these components as
${\Psi}_{\mathrm{d}}(\bm{\xi},t)$ $=\ell_{0}^{-1}\left(\partial_{1}{x}+\partial_{2}{y}\right)-2,$ (16a)
${\Psi}_{\mathrm{r}}(\bm{\xi},t)$ $=\ell_{0}^{-1}\left(\partial_{1}{y}-\partial_{2}{x}\right),$ (16b)
using abbreviation $\partial_{\alpha}=\partial/\partial{\xi_{\alpha}}$ ($\alpha\in\{1,2\}$).
Expanding ${\Psi}_{\mathrm{d}}(\bm{\xi},t)$ and ${\Psi}_{\mathrm{r}}(\bm{\xi},t)$ into Fourier series as
$\Psi_{a}(\bm{\xi},t)=\sum_{\mathbf{k}}\psi_{a}(\mathbf{k},t)e^{-\mathrm{i}\mathbf{k}\cdot\bm{\xi}}\quad(a\in\{\mathrm{d},\mathrm{r}\}),$ (17)
we define correlations of the Fourier modes by
$C_{a}(\mathbf{k},t,s)=\mathcal{N}\left\langle{{\psi_{a}(\mathbf{k},t)}{\psi_{a}(-\mathbf{k},s)}}\right\rangle,$ (18)
where $\mathcal{N}$ is a normalizing constant; here we choose $\mathcal{N}$ equal to the total particle number $N$ in the domain.
We expect that the correlations of the deformation gradient field, namely ${C}_{\mathrm{d}}$ and ${C}_{\mathrm{r}}$, carry information about stresses associated with their respective modes of deformation. These stresses can be related to the elastic moduli as was discussed by Klix et al. [4], and they can be viscoelastic or simply viscous. The problem is how this information of rheology is reflected by the behavior of DC and SC.
The displacement of the fluid element labeled by $\bm{\xi}$ is given by
$\mathbf{R}(\bm{\xi},t,s)=\mathbf{r}(\bm{\xi},t)-\mathbf{r}(\bm{\xi},s)$ (19)
on the basis of the mapping in Eq. (15). Using Eq. (19), we can derive a formula for calculating the DC tensor from ${C}_{\mathrm{d}}$ and ${C}_{\mathrm{r}}$ [31, 20]. The formula will be given later as Eqs. (24).
In relating the SC with ${C}_{\mathrm{d}}$ and ${C}_{\mathrm{r}}$, some caution is needed to avoid confusion between $\gamma$ and $\partial_{2}{x}+\partial_{1}{y}$; the former, $\gamma=F_{12}+F_{21}$, is originally a two-time quantity, as it comes from $\mathsf{F}$ in Eq. (10) which depends both on $s$ and on $t$, while the latter is a one-time quantity coming from $\partial\mathbf{r}(\bm{\xi},t)/\partial\bm{\xi}$. More suitably, we notice that $\mathsf{F}$ is the differential quotient of the composite mapping
$\mathbf{r}(s)\mapsto\bm{\xi}=\bm{\xi}(\mathbf{r}(s),s)\mapsto\mathbf{r}(t)=\mathbf{r}(\bm{\xi},t)$ (20)
in which $\mathbf{r}(t)$ and $\mathbf{r}(s)$ are connected by way of $\bm{\xi}$ as a parameter. Denoting the components of the relative deformation gradient tensor with $F_{\alpha\beta}=\partial{r_{\alpha}}(t)/\partial{r_{\beta}}(s)$, we have $\mathrm{d}{r_{\alpha}}(t)=F_{\alpha\beta}\mathrm{d}{r_{\beta}}(s)$, so that Eq. (10) can be written, with $\bm{\xi}$ taken as the independent variable, as
$\frac{\partial\mathbf{r}(\bm{\xi},t)}{\partial\bm{\xi}}=\mathsf{F}\cdot\frac{\partial\mathbf{r}(\bm{\xi},s)}{\partial\bm{\xi}}$ (21a)
or, in terms of components,
$\begin{bmatrix}\partial_{1}{x}&\partial_{2}{x}\\ \partial_{1}{y}&\partial_{2}{y}\end{bmatrix}_{t}=\begin{bmatrix}F_{11}&F_{12}\cr F_{21}&F_{22}\end{bmatrix}\begin{bmatrix}\partial_{1}{x}&\partial_{2}{x}\\ \partial_{1}{y}&\partial_{2}{y}\end{bmatrix}_{s},$ (21b)
with subscripted $t$ or $s$ denoting the time argument. Note that $\partial\mathbf{r}/\partial\bm{\xi}$ in Eq. (21b) is connected to ${\Psi}_{\mathrm{d}}$ and ${\Psi}_{\mathrm{r}}$ through Eq. (16). With this connection taken into account, Eq. (21b) provides a foundation for calculation of the SC, as $\gamma=F_{12}+F_{21}$.
III.2 Formulae between the correlations under consideration
On the basis of the label variable formulation of the displacement field $\mathbf{R}(\bm{\xi},t,s)$ in the previous subsection, here we interrelate three kinds of correlations: the DC, the SC, and the correlations of the deformation gradients (namely ${C}_{\mathrm{d}}$ and ${C}_{\mathrm{r}}$). We start by reviewing a formula connecting ${C}_{\mathrm{d}}$ and ${C}_{\mathrm{r}}$ to the DC [27, 20]. Subsequently, we perform analogous calculations for the SC, reproducing the formula reported by Illing et al. [19]. Using relations obtained in course of these calculations, we derive a relation interconnecting the DC and the SC.
III.2.1 Deformation gradient correlations to DC
Let us begin with a formula to calculate the DC tensor, $\mathsf{X}=\left\langle{\mathbf{R}\otimes\mathbf{R}}\right\rangle_{\tilde{\mathbf{d}}}$, from the deformation gradient correlations, ${C}_{\mathrm{d}}$ and ${C}_{\mathrm{r}}$. The starting point is the displacement field $\mathbf{R}(\bm{\xi},t,s)$ in Eq. (19). On the understanding that the initial separation $\tilde{\mathbf{d}}$ in the physical space is equivalent to the label-space separation $\bm{\xi}_{*}=\tilde{\mathbf{d}}/\ell_{0}$ on average, we calculate the DC as
$\mathsf{X}(\tilde{\mathbf{d}},t,s)$ $\simeq\left\langle{\mathbf{R}(\bm{\xi},t,s)\otimes\mathbf{R}(\bm{\xi}^{\prime},t,s)}\right\rangle_{\bm{\xi}-\bm{\xi}^{\prime}=\tilde{\mathbf{d}}/\ell_{0}}$
$=\left\langle{\mathbf{R}(\bm{\xi}^{\prime}+\bm{\xi}_{*},t,s)\otimes\mathbf{R}(\bm{\xi}^{\prime},t,s)}\right\rangle.$ (22)
Note that the expression on the right side of Eq. (22) is actually independent of $\bm{\xi}^{\prime}$ due to the space-translational symmetry.
By expressing the displacement field $\mathbf{R}(\bm{\xi},t,s)$ in terms of the Fourier modes of the deformation gradient tensor, we can derive a formula that relates the DC tensor to the correlations of $\psi_{a}$, where $a\in\{\mathrm{d},\mathrm{r}\}$ [27, 20]. We refer to it as the Alexander–Pincus formula, naming it after the authors of Ref. [31]. In writing this formula, it is convenient to introduce $C_{a}^{0}(\mathbf{k},s)=C_{a}(\mathbf{k},s,s)$ and
$C^{\Delta}_{a}(\mathbf{k},t,s)$ $=\frac{C_{a}^{0}(\mathbf{k},s)+C_{a}^{0}(\mathbf{k},t)}{2}-C_{a}(\mathbf{k},t,s)$
$={\frac{N}{2}}\left\langle{\left|{{\psi_{a}(\mathbf{k},t)}-{\psi_{a}(\mathbf{k},s)}}\right|^{2}}\right\rangle.$ (23)
The Alexander–Pincus formula then reads [27, 20]
$\mathsf{X}(\ell_{0}\bm{\xi}_{*},t,s)=\frac{\ell_{0}^{2}}{2\pi^{2}}\tensor{I}_{\mathrm{d}}+\frac{\ell_{0}^{2}}{2\pi^{2}}\tensor{I}_{\mathrm{r}},$ (24a)
where
$\tensor{I}_{\mathrm{d}}$ $=\iint{C}_{\mathrm{d}}^{\Delta}(\mathbf{k},t,s)\begin{bmatrix}k_{1}^{2}&{k_{1}}{k_{2}}\\ {k_{2}}{k_{1}}&k_{2}^{2}\end{bmatrix}\frac{e^{-\mathrm{i}\mathbf{k}\cdot\bm{\xi}_{*}}}{\mathbf{k}^{4}}\mathrm{d}{k_{1}}\mathrm{d}{k_{2}},$ (24b)
$\tensor{I}_{\mathrm{r}}$ $=\iint{C}_{\mathrm{r}}^{\Delta}(\mathbf{k},t,s)\begin{bmatrix}k_{2}^{2}&\!{-{k_{1}}{k_{2}}}\\ -{k_{2}}{k_{1}}&k_{1}^{2}\end{bmatrix}\frac{e^{-\mathrm{i}\mathbf{k}\cdot\bm{\xi}_{*}}}{\mathbf{k}^{4}}\mathrm{d}{k_{1}}\mathrm{d}{k_{2}}.$ (24c)
The assumption of statistical isotropy implies $C_{a}^{\Delta}(\mathbf{k},t,s)=C_{a}^{\Delta}(k,t,s)$ and suggests to introduce polar coordinates in the $\mathbf{k}$-space and the $\bm{\xi}_{*}$-space,
$\mathbf{k}=\begin{bmatrix}k_{1}\\ k_{2}\end{bmatrix}=k\begin{bmatrix}\cos\varphi\\ \sin\varphi\end{bmatrix},\quad\bm{\xi}_{*}=\begin{bmatrix}\xi_{*1}\\ \xi_{*2}\end{bmatrix}=\xi_{*}\begin{bmatrix}\cos\varphi_{*}\\ \sin\varphi_{*}\end{bmatrix},$ (25)
so that we can rewrite the integrals in Eqs. (24) as
$\tensor{I}_{\mathrm{d}}$ $=\iint{C}_{\mathrm{d}}^{\Delta}(k,t,s)\begin{bmatrix}\cos^{2}\varphi&\cos\varphi\sin\varphi\\ \,\sin\varphi\cos\varphi&\sin^{2}\varphi\end{bmatrix}\frac{e^{-\mathrm{i}{k}\xi_{*}\cos(\varphi-\varphi_{*})}}{k}\,\mathrm{d}{k}\,\mathrm{d}{\varphi}$
$=\pi\int_{0}^{\infty}{C}_{\mathrm{d}}^{\Delta}(k,t,s)\left\{\begin{bmatrix}1&0\\ 0&1\end{bmatrix}J_{0}(k\xi_{*})-\begin{bmatrix}\cos{2\varphi_{*}}&\,\sin{2\varphi_{*}}\\ \sin{2\varphi_{*}}&-\cos{2\varphi_{*}}\end{bmatrix}J_{2}(k\xi_{*})\right\}\frac{\mathrm{d}{k}}{k},$ (26a)
$\tensor{I}_{\mathrm{r}}$ $=\iint{C}_{\mathrm{r}}^{\Delta}(k,t,s)\begin{bmatrix}\sin^{2}\varphi&\!{-\cos\varphi\sin\varphi}\\ -\sin\varphi\cos\varphi&\cos^{2}\varphi\end{bmatrix}\frac{e^{-\mathrm{i}{k}\xi_{*}\cos(\varphi-\varphi_{*})}}{k}\,\mathrm{d}{k}\,\mathrm{d}{\varphi}$
$=\pi\int_{0}^{\infty}{C}_{\mathrm{r}}^{\Delta}(k,t,s)\left\{\begin{bmatrix}1&0\\ 0&1\end{bmatrix}J_{0}(k\xi_{*})+\begin{bmatrix}\cos{2\varphi_{*}}&\,\sin{2\varphi_{*}}\\ \sin{2\varphi_{*}}&-\cos{2\varphi_{*}}\end{bmatrix}J_{2}(k\xi_{*})\right\}\frac{\mathrm{d}{k}}{k},$ (26b)
with the Bessel functions, $J_{0}$ and $J_{2}$, arising from integration over $\varphi$ [32]. This form can be more useful than Eqs. (24) in evaluating the integrals when ${C}_{\mathrm{d}}(k,t,s)$ and ${C}_{\mathrm{r}}(k,t,s)$ are known.
III.2.2 Deformation gradient correlations to SC
In parallel to the Alexander–Pincus formula that expresses the DC in terms of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$, it is possible to derive an analogous formula for the SC. The derivation starts with solving Eq. (21b) for $\mathsf{F}$, which gives
$\mathsf{F}=\begin{bmatrix}F_{11}&F_{12}\\ F_{21}&F_{22}\end{bmatrix}=\begin{bmatrix}\partial_{1}{x}&\partial_{2}{x}\\ \partial_{1}{y}&\partial_{2}{y}\end{bmatrix}_{t}\left(\rho\begin{bmatrix}~{}\;\partial_{2}{y}&\!{-\partial_{2}{x}}\\ -\partial_{1}{y}&\!{~{}\;\partial_{1}{x}}\end{bmatrix}\right)_{s}$ (27)
with the Jacobian determinant $\partial(x,y)/\partial(\xi_{1},\xi_{2})=1/\rho$ taken into account. The time arguments are subscripted as in Eqs. (21b). The off-diagonal components of $\mathsf{F}$ are then evaluated:
$F_{12}$ $=\rho_{s}\left(-\left.{\frac{\partial x}{\partial\xi_{1}}}\right|_{t}\left.{\frac{\partial x}{\partial\xi_{2}}}\right|_{s}+\left.{\frac{\partial x}{\partial\xi_{2}}}\right|_{t}\left.{\frac{\partial x}{\partial\xi_{1}}}\right|_{s}\,\right)$
$\quad{}\simeq\ell_{0}^{-1}\left(\partial_{2}{x(\bm{\xi},t)}-\partial_{2}{x(\bm{\xi},s)}\right),$ (28a)
$F_{21}$ $=\rho_{s}\left(~{}\left.{\frac{\partial y}{\partial\xi_{1}}}\right|_{t}\left.{\frac{\partial y}{\partial\xi_{2}}}\right|_{s}-\left.{\frac{\partial y}{\partial\xi_{2}}}\right|_{t}\left.{\frac{\partial y}{\partial\xi_{1}}}\right|_{s}\,\right)$
$\quad{}\simeq\ell_{0}^{-1}\left(\partial_{1}{y(\bm{\xi},t)}-\partial_{1}{y(\bm{\xi},s)}\right),$ (28b)
where the lowest-order approximation for the diagonal components, $\partial_{1}{x}\simeq\partial_{2}{y}\simeq\ell_{0}$, is taken into account [33]. Upon substitution into the definition of $\gamma$, namely Eq. (11), we find
$\gamma=F_{12}+F_{21}\\ {}\simeq\ell_{0}^{-1}\left({}\partial_{2}{x(\bm{\xi},t)}-\partial_{2}{x(\bm{\xi},s)}+\partial_{1}{y(\bm{\xi},t)}-\partial_{1}{y(\bm{\xi},s)}\right),$ (29)
which is simplified by using the displacement field, $\mathbf{R}=\mathbf{r}(\bm{\xi},t)-\mathbf{r}(\bm{\xi},s)$, as
$\gamma=\ell_{0}^{-1}\left(\partial_{1}{R_{y}}+\partial_{2}{R_{x}}\right).$ (30)
The expression of $\gamma$ in Eq. (30) makes it possible to calculate the SC by evaluating
$\chi_{\gamma}=\chi_{\gamma}(\ell_{0}\bm{\xi}_{*},t,s)=\left\langle{\gamma(\bm{\xi}^{\prime}+\bm{\xi}_{*},t,s)\gamma(\bm{\xi}^{\prime},s,t)}\right\rangle,$ (31)
which is supposed to result in an integral form analogous to Eq. (24) for DC. In fact, it reproduces a formula briefly stated by Illing et al. [19], which reads
$\chi_{\gamma}$ $=\frac{1}{{2\pi^{2}}}\iint{C}_{\mathrm{d}}^{\Delta}(\mathbf{k},t,s)\frac{4{k_{1}^{2}}{k_{2}^{2}}}{\mathbf{k}^{4}}e^{-\mathrm{i}\mathbf{k}\cdot\bm{\xi}_{*}}\mathrm{d}{k_{1}}\mathrm{d}{k_{2}}$
$\quad{}+\frac{1}{{2\pi^{2}}}\iint{C}_{\mathrm{r}}^{\Delta}(\mathbf{k},t,s)\frac{(k_{1}^{2}-k_{2}^{2})^{2}}{\mathbf{k}^{4}}e^{-\mathrm{i}\mathbf{k}\cdot\bm{\xi}_{*}}\mathrm{d}{k_{1}}\mathrm{d}{k_{2}}$ (32)
in our notation.
To evaluate Eq. (31), we operate Eq. (30) with ${\partial_{1}}^{2}+{\partial_{2}}^{2}$ and rearrange the terms so as to express the result in terms of ${\Psi}_{\mathrm{d}}$ and ${\Psi}_{\mathrm{r}}$. Taking notice of the relations
$\partial_{1}{R_{x}}+\partial_{2}{R_{y}}={\Psi}_{\mathrm{d}}(\bm{\xi},t)-{\Psi}_{\mathrm{d}}(\bm{\xi},s),$ (33a)
$\partial_{1}{R_{y}}-\partial_{2}{R_{x}}={\Psi}_{\mathrm{r}}(\bm{\xi},t)-{\Psi}_{\mathrm{r}}(\bm{\xi},s),$ (33b)
we find
$\left({\partial_{1}}^{2}+{\partial_{2}}^{2}\right)\gamma$ $=\ell_{0}^{-1}\left[2\partial_{1}\partial_{2}(\partial_{1}{R_{x}}+\partial_{2}{R_{y}})+(\partial_{1}^{2}-\partial_{2}^{2})(\partial_{1}{R_{y}}-\partial_{2}{R_{x}})\right]$
$=\ell_{0}^{-1}\left\{2\partial_{1}\partial_{2}{}[{\Psi}_{\mathrm{d}}(\bm{\xi},t)-{\Psi}_{\mathrm{d}}(\bm{\xi},s)]+(\partial_{1}^{2}-\partial_{2}^{2})[{\Psi}_{\mathrm{r}}(\bm{\xi},t)-{\Psi}_{\mathrm{r}}(\bm{\xi},s)]\right\},$ (34)
which is rewritten in Fourier representation as
$\gamma(\bm{\xi},t,s)=\sum_{\mathbf{k}}\frac{e^{-\mathrm{i}\mathbf{k}\cdot\bm{\xi}}}{\mathbf{k}^{2}}\left\{2{k_{1}}{k_{2}}\left[{\psi}_{\mathrm{d}}(\mathbf{k},t)-{\psi}_{\mathrm{d}}(\mathbf{k},s)\right]+\left(k_{1}^{2}-k_{2}^{2}\right)\left[{\psi}_{\mathrm{r}}(\mathbf{k},t)-{\psi}_{\mathrm{r}}(\mathbf{k},s)\right]\right\}.$ (35)
Substituting Eq. (35) into Eq. (31), assuming that the correlations between different modes are negligible (by the same reason as is discussed in derivation of the Alexander–Pincus formula [27]), and taking the continuum limit, we arrive at Eq. (32).
The four-fold symmetry of $\chi_{\gamma}$ is readily shown by rewriting Eq. (32) in polar coordinates. Using the angular variables $\varphi$ and $\varphi_{*}$ in Eq. (25), we have
$\frac{4{k_{1}^{2}}{k_{2}^{2}}}{\mathbf{k}^{4}}=\sin^{2}{2\varphi}=\frac{1-\cos{4\varphi}}{2},$
$\frac{(k_{1}^{2}-k_{2}^{2})^{2}}{\mathbf{k}^{4}}=\cos^{2}{2\varphi}=\frac{1+\cos{4\varphi}}{2},$
$e^{-\mathrm{i}\mathbf{k}\cdot\bm{\xi}_{*}}=e^{-\mathrm{i}{k}\xi_{*}\cos(\varphi-\varphi_{*})}$
for the $\varphi$-dependent factors in the integrands. Subsequently, changing the variable of integration from $\varphi$ to $\varphi-\varphi_{*}$, we find the $\varphi_{*}$-dependence of $\chi_{\gamma}$ to be of the form
$\chi_{\gamma}=\chi_{\gamma}^{(0)}+\chi_{\gamma}^{(4)}\cos{4\varphi_{*}}$ (36)
where
$\chi_{\gamma}^{(0)}$ $={\frac{1}{2\pi}}\int_{0}^{\infty}\left[{C}_{\mathrm{d}}^{\Delta}(\mathbf{k},t,s)+{C}_{\mathrm{r}}^{\Delta}(\mathbf{k},t,s)\right]J_{0}(k\xi_{*})k\mathrm{d}{k},$ (37a)
$\chi_{\gamma}^{(4)}$ $={\frac{1}{2\pi}}\int_{0}^{\infty}\left[-{C}_{\mathrm{d}}^{\Delta}(\mathbf{k},t,s)+{C}_{\mathrm{r}}^{\Delta}(\mathbf{k},t,s)\right]J_{4}(k\xi_{*})k\mathrm{d}{k}$ (37b)
are independent of $\varphi_{*}$. This is almost purely a consequence of the definitions of SC, independent of the behavior of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$. To be more precise, we have assumed only the existence of the label-based displacement field and the isotropy of the system.
III.2.3 Relation between DC and SC
In addition to Eq. (24) for DC and Eq. (32) for SC, now we derive a third relation expressing the SC in terms of $X_{\parallel}$ and $X_{\perp}$. Such a relation must be available in principle, because the Alexander–Pincus formula is invertible in the sense that ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$ can be expressed in terms of the DC [20], and substitution into Eq. (32) then yields $\chi_{\gamma}$.
To derive the relation more straightforwardly, we begin with substituting Eq. (30), expressing $\gamma$ in terms of $\mathbf{R}(\bm{\xi},t,s)$, directly into Eq. (31) that gives $\chi_{\gamma}$. Subsequently, taking the space-translational symmetry into account, we rewrite the expression for SC as
$\chi_{\gamma}$ $=\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})][\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime})]}\right\rangle$
$=-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{1^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{y}}(\bm{\xi}^{\prime})}\right\rangle$
$\qquad{}-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{2^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{x}}(\bm{\xi}^{\prime})}\right\rangle$ (38)
where $\partial_{\alpha^{\prime}}$ stands for $\partial/\partial{\xi^{\prime}_{\alpha}}$ ($\alpha\in\{1,2\}$). In Eq. (38), use is made of “integration by parts”
$\left\langle{\psi\partial_{\alpha^{\prime}}\phi}\right\rangle=-\left\langle{(\partial_{\alpha^{\prime}}\psi)\phi}\right\rangle$
for arbitrary functions $\psi=\psi(\bm{\xi}^{\prime})$ and $\phi=\phi(\bm{\xi}^{\prime})$ such that $\left\langle{\psi\phi}\right\rangle$ is spatially uniform.
Manipulating the differentiations in Eq. (38) with relations such as $\partial_{1^{\prime}}{\mathbf{R}(\bm{\xi}^{\prime}+\bm{\xi}_{*})}=\partial_{1_{*}}{\mathbf{R}(\bm{\xi}^{\prime}+\bm{\xi}_{*})}$ (with the meaning of $\partial_{\alpha_{*}}$ obviously understood), and using the “double-dot” product notation [34], we obtain
$\chi_{\gamma}=-\ell_{0}^{-2}\begin{bmatrix}\partial_{1_{*}}\partial_{1_{*}}&\partial_{1_{*}}\partial_{2_{*}}\\ \partial_{2_{*}}\partial_{1_{*}}&\partial_{2_{*}}\partial_{2_{*}}\end{bmatrix}:\begin{bmatrix}\left\langle{{R_{y}}{R_{y}}}\right\rangle&\left\langle{{R_{x}}{R_{y}}}\right\rangle\\ \left\langle{{R_{y}}{R_{x}}}\right\rangle&\left\langle{{R_{x}}{R_{x}}}\right\rangle\end{bmatrix}$ (39)
where $\left\langle{{R_{x}}{R_{x}}}\right\rangle=\left\langle{{R_{x}(\bm{\xi}^{\prime}+\bm{\xi}_{*})}{R_{x}(\bm{\xi}^{\prime})}}\right\rangle$ etc. Note that the rightmost factor in Eq. (39) is not the DC tensor itself but its rearrangement, with the diagonal components exchanged.
Here we recall Eq. (6) to decompose the DC tensor into the longitudinal and transverse correlations, which reads as
$\mathsf{X}$ $=\begin{bmatrix}\left\langle{{R_{x}}{R_{x}}}\right\rangle&\left\langle{{R_{x}}{R_{y}}}\right\rangle\\ \left\langle{{R_{y}}{R_{x}}}\right\rangle&\left\langle{{R_{y}}{R_{y}}}\right\rangle\end{bmatrix}$
$=X_{\parallel}\begin{bmatrix}{{\cos^{2}}{\varphi_{*}}}\!&\!{\sin{\varphi_{*}}\cos{\varphi_{*}}}\\ {\cos{\varphi_{*}}\sin{\varphi_{*}}}\!&\!{{\sin^{2}}{\varphi_{*}}}\end{bmatrix}$
[…]
$=\frac{X_{\parallel}+X_{\perp}}{2}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}+\frac{X_{\parallel}-X_{\perp}}{2}\begin{bmatrix}\cos{2\varphi_{*}}&{~{}}\sin{2\varphi_{*}}\\ \sin{2\varphi_{*}}&{-\cos{2\varphi_{*}}}\end{bmatrix}$ (40)
in the polar coordinate system for the $\bm{\xi}_{*}$-space. Taking notice of the last line in Eq. (40), we define
$f_{\pm}=X_{\parallel}\pm X_{\perp}$ (41)
for later convenience. Then, with Eq. (39) in mind, we rearrange the matrix components in Eq. (40) as
$\begin{bmatrix}\left\langle{{R_{y}}{R_{y}}}\right\rangle&\left\langle{{R_{x}}{R_{y}}}\right\rangle\\ \left\langle{{R_{y}}{R_{x}}}\right\rangle&\left\langle{{R_{x}}{R_{x}}}\right\rangle\end{bmatrix}\\ {}={\frac{1}{2}}{f_{+}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}+{\frac{1}{2}}{f_{-}}\begin{bmatrix}{-\cos{2\varphi_{*}}}&\sin{2\varphi_{*}}\\ {~{}\sin{2\varphi_{*}}}&\cos{2\varphi_{*}}\end{bmatrix},$ (42)
from which we obtain
$\nabla_{*}\cdot\begin{bmatrix}\left\langle{{R_{y}}{R_{y}}}\right\rangle&\left\langle{{R_{x}}{R_{y}}}\right\rangle\\ \left\langle{{R_{y}}{R_{x}}}\right\rangle&\left\langle{{R_{x}}{R_{x}}}\right\rangle\end{bmatrix}\\ {}=\frac{f_{+}^{\prime}}{2\ell_{0}}\begin{bmatrix}\cos{\varphi_{*}}\\ \sin{\varphi_{*}}\end{bmatrix}+\ell_{0}^{-1}\left(-\frac{f_{-}^{\prime}}{2}+\frac{f_{-}}{\xi}\right)\begin{bmatrix}~{}\cos{3\varphi_{*}}\\ -\sin{3\varphi_{*}}\end{bmatrix},$ (43)
where
$\nabla_{*}$ $=\ell_{0}^{-1}\begin{bmatrix}\partial_{1_{*}}\\ \partial_{2_{*}}\end{bmatrix}$
$=\ell_{0}^{-1}\begin{bmatrix}\cos{\varphi_{*}}\\ \sin{\varphi_{*}}\end{bmatrix}\frac{\partial}{\partial\xi_{*}}+(\ell_{0}\xi_{*})^{-1}\begin{bmatrix}-\sin{\varphi_{*}}\\ ~{}\cos{\varphi_{*}}\end{bmatrix}\frac{\partial}{\partial\varphi_{*}}$
and $f_{\pm}^{\prime}=\partial{f_{\pm}}/\partial{\xi_{*}}$. Finally, calculating the ($\xi_{*}$-space) divergence of Eq. (43), we arrive at
$\chi_{\gamma}=-{\frac{1}{2}}{\ell_{0}^{-2}}{\xi_{*}^{-1}}\frac{\partial}{\partial\xi_{*}}\left(\,\xi_{*}f_{+}^{\prime}\right)\\ {}+\ell_{0}^{-2}\left({\frac{1}{2}}f_{-}^{\prime\prime}-\frac{5}{2}{\xi_{*}}^{-1}f_{-}^{\prime}+4\,{\xi_{*}}^{-2}f_{-}\right)\cos{4\varphi_{*}}.$ (44)
The $\varphi_{*}$-dependence of the expression on the right-hand side Eq. (44) is consistent with Eq. (36). We note that the isotropic part is given in terms of $f_{+}$, while the anisotropic part proportional to $\cos{4\varphi_{*}}$ is determined by $f_{-}$. By comparing Eq. (44) with Eq. (36) and introducing linear operators of Euler–Cauchy type,
$\hat{L}_{+}$ $=-{\frac{1}{2}}\ell_{0}^{-2}\left(\frac{\partial^{2}}{\partial\xi_{*}^{2}}+\xi_{*}^{-1}\frac{\partial}{\partial\xi_{*}}\right),$ (45a)
$\hat{L}_{-}$ $=\ell_{0}^{-2}\left({\frac{1}{2}}\frac{\partial^{2}}{\partial\xi_{*}^{2}}-\frac{5}{2}{\xi_{*}}^{-1}\frac{\partial}{\partial\xi_{*}}+4\,{\xi_{*}}^{-2}\right),$ (45b)
we obtain
$\chi_{\gamma}^{(0)}={\hat{L}_{+}f_{+}},\quad\chi_{\gamma}^{(4)}={\hat{L}_{-}f_{-}},$ (45c)
and Eq. (44) is rewritten more concisely as $\chi_{\gamma}={\hat{L}_{+}f_{+}}+(\hat{L}_{-}f_{-})\cos{4\varphi_{*}}$.
Figure 3: Decomposition of the DC tensor into $f_{\pm}=X_{\parallel}\pm X_{\perp}$ according to Eq. (40); see Fig. 1 (b).
In comparison to Eq. (36) for $\chi_{\gamma}$, which is understood simply as a special case of Fourier decomposition with regard to $\varphi_{*}$, the decomposition of the DC tensor into $f_{+}$ and $f_{-}$ in the form of Eq. (40) may seem more difficult to conceive. Pictorially, this decomposition can be illustrated as in Fig. 3, to be compared with Fig. 1. The component given by $f_{+}=X_{\parallel}+X_{\perp}$ represents correlated motion in which all the neighboring particles are dragged in the same direction, while the so-called backflow pattern is represented by $f_{-}=X_{\parallel}-X_{\perp}$ as a vortex dipole.
III.3 Possible nonequivalence between DC and SC
With the relation in Eqs. (45c) we can calculate the SC from the DC, but this does not necessarily mean that SC and DC are equivalent. The problem is whether the relation is invertible, making it possible to obtain the DC from the SC. Intuitively speaking, since differentiation is involved in the operators $\hat{L}_{\pm}$, some information in DC is likely to be missing from SC.
The invertibility of the linear operators $\hat{L}_{\pm}$ in Eq. (45c) can be discussed by checking their null spaces. The null space of $\hat{L}_{+}$, also known as the kernel of the linear operator [35], is the set of all the solutions $h_{+}$ to the homogeneous equation
$\hat{L}_{+}h_{+}=0,$ (46)
with the domain of the function taken as some physically appropriate range of $\xi_{*}$. Noticing that $\hat{L}_{+}$ is essentially the radial part of the Laplacian operator in the $\bm{\xi}_{*}$-space, we can readily find $h_{+}=A+B\log\xi_{*}$ (with $A$ and $B$ denoting arbitrary constants); in other words, the null space is spanned by $\{1,\log\xi_{*}\}$. If $f_{+}\sim h_{+}$, it is mapped to zero by $\hat{L}_{+}$ in Eqs. (45c) and therefore the information of $f_{+}$ is lost from $\chi_{\gamma}^{(0)}$. This makes the DC and the SC nonequivalent.
Analogously, the general solution to the linear homogeneous equation
$\hat{L}_{-}h_{-}=0$ (47)
comprises the null space of $\hat{L}_{-}$, which is spanned by $\{\xi_{*}^{2},\xi_{*}^{4}\}$. If $f_{-}$ happens to fall into the null space of $\hat{L}_{-}$, the information of $f_{-}$ drops out of $\chi_{\gamma}^{(4)}$.
Thus we have derived analytical relations involving the SCs, which allows us to discuss the possibility that the DCs and the SCs are not equivalent due to the null space of $\hat{L}$. Now let us proceed to numerical study of these analytical results, beginning with specification of the particle system in the next section.
IV Specification of the particle system
Let us specify the particle system as a model liquid, for which we calculate the DC and the SC numerically. We consider a system consisting of $N$ Brownian particles (disks) in a 2D periodic box of the size $L^{2}$, which is basically the same system as in our previous work [20]. The position vectors of the particles, denoted by $\mathbf{r}_{i}$ with $i=1,2,\ldots,N$, are governed by the overdamped Langevin equation
$\mu\dot{\mathbf{r}}_{i}=-\frac{\partial}{\partial\mathbf{r}_{i}}\sum_{j<k}V_{jk}+\mu\mathbf{f}_{i}(t),$ (48)
where $\mu$ is the drag coefficient, and $\mu\mathbf{f}_{i}(t)$ is the thermal fluctuation term with the temperature $T$, corresponding to the bare diffusivity $D={k_{\text{B}}}T/\mu$ and prescribed as a Gaussian random forcing with zero mean and the variance
$𝟙\left\langle{\mathbf{f}_{i}(t)\otimes\mathbf{f}_{j}(t^{\prime})}\right\rangle=2D\delta_{ij}\delta(t-t^{\prime})\openone.$
As the interaction potential $V_{jk}$ between the $j$-th and $k$-th particles (separated by the relative position vector $\mathbf{r}_{ij}$), we adopt the harmonic repulsive potential,
$V_{jk}=\begin{cases}V_{\text{max}}\left(1-\dfrac{\left|{\mathbf{r}_{ij}}\right|}{\sigma}\right)^{2}&(\left|{\mathbf{r}_{ij}}\right|<\sigma)\cr 0&(\text{otherwise})\end{cases}$
with very large barrier $V_{\text{max}}$, nearly equivalent to the hardcore interaction with diameter $\sigma$ (the particles are monodisperse). The inertia is completely ignored.
We prepare the system to be in a statistically steady and homogeneous state in equilibrium at the temperature $T$. The mean density is $\rho_{0}=N/L^{2}$, with which we define $\ell_{0}=1/\sqrt{\rho_{0}}$ as a length scale that represents the typical interparticle distance. The area fraction is $\phi_{\text{area}}=(\pi/4)\sigma^{2}\rho_{0}=(\pi/4)(\sigma/\ell_{0})^{2}$.
V Numerical results
We performed numerical simulation of the system of Brownian disks specified in Sec. IV. The system contains $N=4000$ particles, and $L$ is adjusted so as to give three values of the area fraction: $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$. To mimic the hardcore interaction, we chose $V_{\text{max}}=50$ for $\phi_{\text{area}}=0.50$, $V_{\text{max}}=500$ for $\phi_{\text{area}}=0.60$, and $V_{\text{max}}=5000$ for $\phi_{\text{area}}=0.70$ in units of ${k_{\text{B}}}T$; the time step for numerical integration, $\Delta{t}$, is chosen so as to satisfy $(\Delta{t}/\tau_{0})({k_{\text{B}}}T/V_{\text{max}})=0.1$ where $\tau_{0}=\sigma^{2}/D$.
In what follows, reference scales for nondimensionalization, such as $\sigma$ and $\ell_{0}$, will be shown explicitly as a rule. In referring to the time interval $t_{\Delta}$, however, we will make an exception: we will write simply $t=20$, for example, instead of $t_{\Delta}/\tau_{0}=20$.
V.1 Detectability of DC and SC in only slightly glassy liquids
Table 1: Longwave limiting values of the static structure factor, $S$, and the structural relaxation time, $\tau_{\alpha}$, obtained from 2D particle simulation data. The values of $\ell_{0}/\sigma$ are also included in the table.
$\phi_{\text{area}}$ 0.50 0.60 0.70
$S$ 0.16 0.055 0.025
$\tau_{\alpha}/\tau_{0}$ 0.032 0.050 0.159
$\ell_{0}/\sigma$ 1.253 1.144 1.059
We start with calculating the quantities tabulated in Table 1 for the range of $\phi_{\text{area}}$ under consideration, where $S$ denotes the longwave limiting values of the static structure factor. The $\alpha$ relaxation time, $\tau_{\alpha}$, is determined by the condition that $F_{\text{S}}(k_{0},\tau_{\alpha})=1/e$, with $F_{\text{S}}$ denoting the self part of the intermediate scattering function,
$F_{\text{S}}(k,t_{\Delta})=\left\langle{{\frac{1}{N}}\sum_{i}\exp\left[{\mathrm{i}\mathbf{k}\cdot\mathbf{R}_{i}(t,s)}\right]}\right\rangle\quad(t=s+t_{\Delta}),$ (49)
and $k_{0}=2\pi/\sigma$ [5, 22]. Plots of $F_{\text{S}}(k_{0},t_{\Delta})$ are shown in Fig. 4 along with the MSD, $\left\langle{\left[{\mathbf{R}(t,s)}\right]^{2}}\right\rangle$. For $t_{\Delta}\gg\tau_{\alpha}$, the MSD grows asymptotically in proportion to $t_{\Delta}$, demonstrating that the separate “steps” in Eq. (2) are uncorrelated.
(a)Refer to caption(b)Refer to caption
Figure 4: Time dependence of well-established statistical quantities based on single particle displacement, computed for three different values of the area fraction ($\phi_{\text{area}}=0.5$, $0.6$, and $0.7$). Note that the axis label $t$ actually stands for $t_{\Delta}/\tau_{0}$. (a) The self part of the intermediate scattering function, $F_{\text{S}}(k,t_{\Delta})$, at $k=k_{0}\;(=2\pi/\sigma)$. (b) The MSD.
Noticing the modesty of $\tau_{\alpha}$ in Table 1, which indicates that the liquid is only slightly glassy, we recall the appreciable presence of DCs even in such cases, as was reported in our previous work [20] and reaffirmed in Fig. 1 of the present article. In regard to the displacement $\mathbf{R}$ in the case of such modest $\tau_{\alpha}$, it should be noted that the time averaging implied in Eq. (1) does not suffice by itself to get rid of the noisiness from displacement-based statistical quantities. Even in such cases, with the aid of ensemble averaging, the statistical procedure in Appendix A reveals appreciable presence of DCs, as we have seen in Fig. 1 of the present article and in Figs. 5, 6, and 8 of Ref. [20], not only for $t=0.50$ but also for $t=15.9$. In other words, displacements are definitely correlated in such an only slightly glassy liquid ($\phi_{\text{area}}=0.5$).
(a)Refer to caption(b)Refer to caption
Figure 5: Comparison among three distinct evaluations of the SC for $t=20$ and $\phi_{\text{area}}=0.5$. The two panels correspond to different directions of $\tilde{\mathbf{d}}$: (a) $\varphi_{*}=0$, and (b) $\varphi_{*}=\pi/4$. The red squares indicate direct computation from the particle simulation data with Eq. (12), while indirect evaluation from the DC by way of Eq. (44) is plotted with blue solid circles, and analytical expression in Eq. (63) with ${\mu_{\text{r}}}=0.1$ is shown with a thin solid line in each panel.
Let us proceed to the calculation of the SC. The angular dependence of SC, involving $\cos{4\varphi_{*}}$ as in Eqs. (36) and (44), is almost self-evident: the isotropy of the system implies that the correlation of $\gamma$, which is a component of a second order tensor, should have the fourfold angular symmetry. Instead of the angular dependence, we should rather focus on the dependence on the distance $\tilde{{d}}$, with the angle $\varphi_{*}$ fixed.
Since the SCs and the DCs are theoretically predicted to be related by Eq. (44), we can expect presence of SC in cases in which DC is detectable. In order to validate this qualitative expectation, along with the quantitative relation in Eq. (44), we computed $\chi_{\gamma}$ from the simulation data, using the definition in Eq. (12) and following the numerical procedure in the supplemental material of Illing et al. [19].
The values of SC thus calculated by direct usage of Eq. (12) are plotted in Fig. 5. For comparison, we have also included the values calculated indirectly by way of Eq. (44) from the DCs. It is evident from Fig. 5 that the values of SC are in reasonable agreement, except for the shorter range with $\tilde{{d}}<7\sigma$ for which it is difficult to obtain reliable values of SC. Thus, at least in the case of Fig. 5 and for $\tilde{{d}}>7\sigma$, we have demonstrated the detectability of the SC as well as the validity of Eq. (44) that relates the DC and the SC.
The solid lines in Fig. 5 represent analytical curves based on an approximate expression for SC, to be given later as Eq. (63) in Subsec. VI.1. The behavior of the curve, decaying to zero as $\tilde{{d}}$ increases, is qualitatively consistent with the numerical results, though some disagreement for small $\tilde{{d}}$ is also visible.
V.2 Indication of nonequivalence between SC and DC
Having seen the validity of Eq. (44) relating the DC and the SC, it is natural to ask whether the two kinds of correlations are equivalent. We start answering this question with comparative study of their dependences on $\phi_{\text{area}}$.
In Fig. 6(a), the SCs (at $\varphi_{*}=0$) for $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$ are compared, while the corresponding plots of $X_{\perp}$, the transverse DC, is shown in Fig. 6(b). Evidently, $\chi_{\gamma}$ in Fig. 6(a) exhibits weaker dependence on $\phi_{\text{area}}$ than $X_{\perp}$ in Fig. 6(b). In other words, DCs are more sensitive to $\phi_{\text{area}}$ than SCs; this difference suggests that DCs and SCs are not equivalent.
(a)Refer to caption(b)Refer to caption
Figure 6: Dependence of the correlations on the area fraction ($\phi_{\text{area}}=0.5$, $0.6$, and $0.7$). (a) SC with $\varphi_{*}=0$. (b) Transverse DC.
By closer observation on Fig. 6, we find a significant difference in the shorter-distance range with $\tilde{{d}}<7\sigma$. As was noticed in the previous subsection, it is difficult to obtain reliable values of SC in the shorter-distance range. Contrastively, the values of DCs seem still reliable in this range, as the numerical values appear to behave consistently throughout the almost entire range of $\tilde{{d}}>\ell_{0}$. We also notice that, in the cases of $\phi_{\text{area}}=0.60$ and $0.70$, there is a significant increase in $\mathsf{X}$ as $\tilde{{d}}$ approaches zero. No counterpart of this $\phi_{\text{area}}$-dependent increase in DC seems to be recognizable in SC.
These observations on the behavior of DC and SC can be explained in terms of the null space of the operator $\hat{L}_{+}$ given in Eq. (45a). As was discussed in Subsec. III.3, the mapping from $f_{\pm}$ to $\chi_{\gamma}$ is not invertible in general, as long as $f_{\pm}$ may fall into the null space of $\hat{L}_{\pm}$. This makes the DC and the SC nonequivalent.
Let us discuss $\hat{L}_{+}$ and $\hat{L}_{-}$ separately. The null space of $\hat{L}_{-}$ consists of the general solution to Eq. (47), spanned by $\{\xi_{*}^{2},\xi_{*}^{4}\}$, whose relevance can be ruled out by the observation that the DCs do not exhibit such a strong divergence for $\xi_{*}=\tilde{{d}}/\ell_{0}\to+\infty$. Contrastively, from the solutions to Eq. (46), we find that $\ln\xi_{*}$ belongs to the null space of $\hat{L}_{+}$, which seems to give a consistent explanation of the numerical observations. If $f_{+}$ behaves as $\ln\xi_{*}$ in some shorter-distance range of $\xi_{*}$, the components of the DC tensor also diverge as
$X_{\parallel}\sim X_{\perp}\sim\ln\xi_{*},$
while the information is lost from $\chi_{\gamma}$, because $\hat{L}_{+}\ln\xi_{*}=0$. Thus the numerical observations indicating nonequivalence between SC and DC can be explained if the DC behaves logarithmically for small $\xi_{*}$.
V.3 Logarithmic behavior of DCs
(a)Refer to caption(b)Refer to caption
Figure 7: DCs plotted against the spatial separation $\tilde{{d}}$, with the $\tilde{{d}}$-axis scaled logarithmically. In both panels, DCs computed at $t=20$ for $\phi_{\text{area}}=0.5$, $0.6$, and $0.7$ are shown. (a) Transverse DCs. (b) Longitudinal DCs.
To establish the explanation of the nonequivalence between SC and DC due to the null space of $\hat{L}_{+}$, now let us demonstrate the presence of logarithmic behavior in DC. This is demonstrated by plotting the components of $\mathsf{X}$ against the logarithm of $\xi_{*}=\tilde{{d}}/\ell_{0}$ and then finding the range of $\xi_{*}$ over which the plots form straight lines.
In Fig. 7 we have such a semilogarithmic plot (of linear–log type). A narrow but recognizable range of straight-line behavior is present. We also notice that the slope of the plots for $X_{\parallel}$ is nearly the same as that for $X_{\perp}$. Besides, the slope becomes steeper as $\phi_{\text{area}}$ is increased.
V.4 Time dependence of SC and DC: diffusive scaling
(a) Refer to caption(b) Refer to caption
Figure 8: SCs versus $\tilde{{d}}$ for three different time interval, computed for (a) $\varphi_{*}=0$ and (b) $\varphi_{*}=\pi/4$.
It seems rather surprising that correlations at finite $\xi_{*}$, such as DC and SC, persists for $t_{\Delta}\gg\tau_{\alpha}$. Some authors proposed to interpret this persistence as an accumulated effect of many events with shorter correlation time; this was proposed by Doliwa and Heuer [11] in regard to DC, and later by Chattoraj and Lemaître [18] for SC. This interpretation leads to the prediction that $\mathsf{X}(\tilde{\mathbf{d}},t_{\Delta})$ and $\chi_{\gamma}(\tilde{\mathbf{d}},t_{\Delta})$ are proportional to $t_{\Delta}$ (with $\tilde{\mathbf{d}}$ fixed), as reviewed at the end of Subsec. II.3.
This prediction, however, is inconsistent with the numerical results shown in Fig. 8, in which plots of SC are compared for three different values of $t_{\Delta}$. The SCs (with $\tilde{\mathbf{d}}$ fixed) do not grow in proportion to $t_{\Delta}$, but rather seem to exhibit some different type of $t_{\Delta}$-dependence through a similarity variable.
(a)Refer to caption(b)Refer to caption
Figure 9: Test for expressibility of SC in terms of the similarity variable $\vartheta=\xi_{*}/(2\sqrt{D^{\text{c}}_{*}{t}})$. The same data as in Fig. 8 are used and shown to collapse into single curves for (a) $\varphi_{*}=0$ and (b) $\varphi_{*}=\pi/4$.
In our previous work [20] on DC (mainly with $\phi_{\text{area}}=0.5$), we found that $X_{\parallel}$ and $X_{\perp}$ are expressible in terms of a similarity variable,
$\vartheta{}=\frac{\tilde{{d}}}{2\sqrt{{D^{\text{c}}}t_{\Delta}\mathstrut}}=\frac{\xi_{*}}{2\sqrt{{D^{\text{c}}_{*}}t_{\Delta}\mathstrut}},$ (50)
including the diffusive length scale $2\sqrt{{D^{\text{c}}}t_{\Delta}}$ where $D^{\text{c}}=D/S$ and $D^{\text{c}}_{*}=D^{\text{c}}/\ell_{0}^{2}$; see Eq. (2.10) and Eq. (4.38) in Ref. [20] about this length scale. The result suggests that the cages are nested to form a self-similar structure in the space–time.
Since the SCs are related with the DCs by Eq. (44), we expect that the SCs are also expressible in terms of the same similarity variable $\vartheta$. This is verified by taking the data of SC in Fig. 8 and replotting them against $\vartheta$. As a result, the plots for different $t_{\Delta}$ are seen to collapse into a single curve, as is shown in Fig. 9. Note that the error bars in Fig. 8 are magnified in Fig. 9 by the factor of $\xi^{2}$ and therefore appears to be large.
(a)Refer to caption(b)Refer to caption
Figure 10: $f_{\pm}$ defined by Eq. (41), computed for $\phi_{\text{area}}=0.60$ and plotted against $\vartheta$.
Figure 11: Replot of $f_{-}$ based on the same data as in Fig. 10(b). The curves collapse onto a single straight line by plotting $\vartheta^{2}f_{-}$ against $\vartheta^{2}$, with the vertical axis in the logarithmic scale.
Finally, let us examine $f_{\pm}$ defined by Eq. (41). They are plotted against the similarity variable $\vartheta$ in Fig. 10, with the $\vartheta$-axis in logarithmic scale. The data of $f_{+}$ for different time intervals are seen to collapse on a single curve in Fig. 10(a), and the curve is close to a straight line (indicating logarithmic behavior) for $\vartheta<0.1$, i.e. for distances much shorter than the diffusive length scale. Thus $f_{+}$ is shown to behave logarithmically at shorter distances. Contrastively, as is seen in Fig. 10(b), the behavior of $f_{-}$ in the range of $\vartheta<0.1$ is not logarithmic at all. For $10.3<\vartheta<$ 1, the data of $f_{-}$ are seen to collapse on a single curve, but the semi-logarithmic plot in Fig. 10 fails to make this curve a straight line. A different kind of plot gives a better explanation for $f_{-}$: as is shown in Fig. 11, the curve is straightened by plotting $\vartheta^{2}f_{-}$ in logarithmic scale against $\vartheta^{2}$ in linear scale. This implies that $f_{-}$ behaves as $\vartheta^{-2}e^{-\vartheta^{2}}$ at longer distances ($\vartheta\sim 1$).
VI Discussion
For the generic framework proposed in Sec. III to relate the SCs with other correlations, we verified numerically, in the previous section, that the SC can be obtained from the components of the DC tensor. We have also shown that the computed DCs behave logarithmically on shorter lengthscales. The slope or the amplitude of the logarithmic part depends on the area fraction $\phi_{\text{area}}$, but the corresponding $\phi_{\text{area}}$-dependence is missing from the SCs, as was confirmed numerically in Sec. V with Fig. 6.
Now let us discuss what kind of information can be read from the logarithmic behavior of the DC. We start with a rough but simple modeling by fluctuating elastic media, which provides specific expressions of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$ as inputs into the Alexander–Pincus formula (24) to calculate the DCs concretely. Subsequently, we will extend our discussion to a wider class of models for $C_{a}^{\Delta}$, showing that the logarithmic regime of the DC can be caused by a certain kind of “caged” behavior of $C_{a}^{\Delta}$.
VI.1 Elastic modeling
The formulae in Subsec. III.2, relating $C_{a}^{\Delta}$ to the DC and the SC, are generic in the sense that dynamics of $\psi_{a}$ are not specified. Although the dynamics for the particles are given by the Langevin equation (48), it is not obvious how the dynamics are projected onto those of $\psi_{a}$ to yield manageable expression of $C_{a}^{\Delta}$. To proceed further, here we assume some approximate dynamics that give
${C}_{\mathrm{d}}^{\Delta}(\mathbf{k},t,s)={C}_{\mathrm{d}}^{\Delta}(k,t_{\Delta})$ $=S\,\left(1-e^{-D^{\text{c}}_{*}\mathbf{k}^{2}t_{\Delta}}\right),$ (51a)
${C}_{\mathrm{r}}^{\Delta}(\mathbf{k},t,s)={C}_{\mathrm{r}}^{\Delta}(k,t_{\Delta})$ $=\frac{S}{{\mu_{\text{r}}}}\,\left(1-e^{-{\mu_{\text{r}}}D^{\text{c}}_{*}\mathbf{k}^{2}t_{\Delta}}\right),$ (51b)
as is proposed in the discussing section of Ref. [20]; here $S$ and ${\mu_{\text{r}}}$ are positive constants [36], and $D^{\text{c}}_{*}=D^{\text{c}}/\ell_{0}^{2}=(D/S)/{\ell_{0}^{2}}$. To be consistent with the known dynamics of $\rho(\mathbf{r},t)$, the constant $S$ is chosen to be equal to the longwave limiting value of the static structure factor (see Table 1).
The correlations proposed in Eq. (51) are understood as a result of elastic modeling for the dynamics of the deformation gradient field, which may be formulated as Langevin equations in the following form:
$(\partial_{t}+D^{\text{c}}_{*}\,\mathbf{k}^{2}){\psi}_{\mathrm{d}}(\mathbf{k},t)$ $=\check{f}_{\text{d}}(\mathbf{k},t),$ (52a)
$(\partial_{t}+{\mu_{\text{r}}}D^{\text{c}}_{*}\,\mathbf{k}^{2}){\psi}_{\mathrm{r}}(\mathbf{k},t)$ $=\check{f}_{\text{r}}(\mathbf{k},t),$ (52b)
where $\check{f}_{\text{d}}$ and $\check{f}_{\text{r}}$ are thermal fluctuation terms with the temperature $T$. The term with the coefficient $D^{\text{c}}_{*}\propto{k_{\text{B}}}T/S$ in Eq. (52a) represents “restoring force” exerted upon ${\psi}_{\mathrm{d}}$ by the elastic medium with the bulk modulus ${k_{\text{B}}}T/S$. We note, parenthetically, that a 1D version of Eq. (52a) has often been used to describe continuum dynamics of a chain of colloidal particles in a channel [37, 27, 38].
In parallel to the term with $D^{\text{c}}_{*}$ in Eq. (52a), the presence of the term with ${\mu_{\text{r}}}D^{\text{c}}_{*}$ in Eq. (52b) means introducing the shear modulus ${\mu_{\text{r}}}{k_{\text{B}}}T/S$ into the model. While this is a natural modeling of DCs in the case of glass solids with plateau modulus or with idealization of $\tau_{\alpha}\to\infty$ [10, 4, 12], it may seem questionable to what extent the elastic modeling is applicable to the case of glassy liquids with modest $\tau_{\alpha}$. Here we regard the elastic modeling as a convenient starting point for searching more sophisticated description of $C_{a}^{\Delta}$ with a wider range of applicability, involving solid-like elasticity at some scale. A recent experiment on confined liquid glyserol [39] may support relevance of such solid-like elasticity in liquids far away from glass transition.
Given the numerical data of the DCs, we focus our attention to the ratio of the coefficients $S$ and $S/{\mu_{\text{r}}}$ in Eq. (51). In principle, ${\mu_{\text{r}}}$ can be computed from the data in the Fourier space as the ratio of the saturation values of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$:
${\mu_{\text{r}}}=\lim_{t_{\Delta}\to\infty}\frac{{C}_{\mathrm{d}}^{\Delta}(k,t_{\Delta})}{{C}_{\mathrm{r}}^{\Delta}(k,t_{\Delta})}.$ (53)
In the case of solidified glass, ${\mu_{\text{r}}}$ represents the ratio of the shear modulus to the bulk modulus, which certainly makes sense for any $k$ small enough to justify elastic modeling and with $t_{\Delta}\to\infty$ understood within the range of $t_{\Delta}\ll\tau_{\alpha}$ [4, 12]. Here we relax this restriction on $t_{\Delta}$, however, expecting that Eq. (51) can be still valid for $t_{\Delta}>\tau_{\alpha}$ in some range of $k$ and may serve as a useful starting point for discussion.
With ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$ given in Eq. (51), we can calculate the DC and the SC analytically, using the formulae in Subsec. III.2.
Let us begin with the Alexander–Pincus formula (24) for the DC tensor. Before staring the calculation, we note that the mathematical procedure is somewhat simplified by utilizing
$\frac{\partial}{\partial t_{\Delta}}\left({{C}_{\mathrm{d}}^{\Delta}},\;{C}_{\mathrm{r}}^{\Delta}\right)=\frac{Dk^{2}}{\ell_{0}^{2}}\,\left({e^{-D^{\text{c}}_{*}{k^{2}}t_{\Delta}}},\;{e^{-{\mu_{\text{r}}}D^{\text{c}}_{*}{k^{2}}t_{\Delta}}}\right).$ (54)
Combining Eq. (54) with the Alexander–Pincus formula in the polar coordinate form in Eq. (26), we have
$\frac{\partial\tensor{I}_{\mathrm{d}}}{\partial t_{\Delta}}=\frac{{\pi}D}{\ell_{0}^{2}}\int_{0}^{\infty}{e^{-D^{\text{c}}_{*}{k^{2}}t_{\Delta}}}\left\{\begin{bmatrix}1&0\\ 0&1\end{bmatrix}J_{0}(k\xi_{*})-\begin{bmatrix}\cos{2\varphi_{*}}&\,\sin{2\varphi_{*}}\\ \sin{2\varphi_{*}}&-\cos{2\varphi_{*}}\end{bmatrix}J_{2}(k\xi_{*})\right\}k\mathrm{d}{k},$ (55a)
$\frac{\partial\tensor{I}_{\mathrm{r}}}{\partial t_{\Delta}}=\frac{{\pi}D}{\ell_{0}^{2}}\int_{0}^{\infty}{e^{-{\mu_{\text{r}}}D^{\text{c}}_{*}{k^{2}}t_{\Delta}}}\left\{\begin{bmatrix}1&0\\ 0&1\end{bmatrix}J_{0}(k\xi_{*})+\begin{bmatrix}\cos{2\varphi_{*}}&\,\sin{2\varphi_{*}}\\ \sin{2\varphi_{*}}&-\cos{2\varphi_{*}}\end{bmatrix}J_{2}(k\xi_{*})\right\}k\mathrm{d}{k}.$ (55b)
By evaluating the wavenumber integrals according to Appendix B and then calculating the antiderivatives with regard to $t_{\Delta}$, we obtain $\tensor{I}_{\mathrm{d}}$ and $\tensor{I}_{\mathrm{r}}$ as functions of $\xi_{*}$, $\varphi_{*}$ and $t_{\Delta}$. The result turns out to be expressible in terms of the similarity variable $\vartheta$ in Eq. (50); then, substituting the result into Eq. (24a) and rearranging the terms by their $\varphi_{*}$-dependence into the form of Eq. (40), we obtain the longitudinal and transverse DCs as functions of $\vartheta$. Using $f_{\pm}$ to express these correlations as
$X_{\parallel}={\frac{1}{2}}\left({f_{+}}+{f_{-}}\right),\quad X_{\perp}={\frac{1}{2}}\left({f_{+}}+{f_{-}}\right)$ (56)
in accordance with Eq. (41), we have
$f_{+}$ $=\frac{S}{2\pi}\ell_{0}^{2}\left[E_{1}(\vartheta^{2})+\frac{E_{1}({\vartheta^{2}}/{\mu_{\text{r}}})}{{\mu_{\text{r}}}}\right],$ (57)
$f_{-}$ $=\frac{S}{2\pi}\ell_{0}^{2}\times\frac{e^{-\vartheta^{2}}-e^{-{\vartheta^{2}}/{\mu_{\text{r}}}}}{\vartheta^{2}},$ (58)
where $E_{1}(~{}\cdot~{})$ denotes the exponential integral defined by [32]
$E_{1}(w)=\int_{w}^{\infty}\frac{\exp(-z)}{z}\mathrm{d}{z}.$ (59)
With Eq. (57) in hand, we can show readily that $f_{+}$ behaves logarithmically on shorter lengthscales. In terms of $\vartheta$ defined in Eq. (50), by the “shorter lengthscales” we mean the range of $\tilde{{d}}=\ell_{0}\xi_{*}$ satisfying both $\vartheta\ll 1$ and $\tilde{{d}}>\ell_{0}$. Using
$E_{1}(w)\simeq-\ln{w}-\gamma_{\text{EM}}+w-\frac{w^{2}}{4}+\cdots$ (60)
for small $w$ (with $\gamma_{\text{EM}}\approx 0.5772$ denoting the Euler–Mascheroni constant), it is straightforward to obtain
$\frac{f_{+}}{(S/2\pi)\ell_{0}^{2}}=-\frac{2(1+{\mu_{\text{r}}})}{{\mu_{\text{r}}}}\ln\vartheta+\mathcal{O}(1)$ (61)
for small $\vartheta$. It is also easy to show
$\frac{f_{-}}{(S/2\pi)\ell_{0}^{2}}=\frac{1-{\mu_{\text{r}}}}{{\mu_{\text{r}}}}+\mathcal{O}(\vartheta^{2})$ (62)
from Eq. (58). The presence of the logarithmic behavior in Eq. (61) and its absence from Eq. (62) are qualitatively consistent with the numerical results in Fig. 10 for small $\vartheta\;(<0.1)$. Besides, with regard to the behavior of $f_{-}$ for large $\vartheta$, we find Eq. (58) to be consistent with our numererical results, as shown in Fig. 11.
The analytical expressions of $f_{\pm}$ in Eqs. (61) and (62) allow us to extract information of elasticity from the numerical data of the particle system, on the assumption that the elastic model is quantitatively valid. They may provide useful alternatives to Eq. (53), as $\mathsf{X}(\tilde{\mathbf{d}},t_{\Delta})$ is easier to compute than its Fourier counterpart used in Eq. (53).
Starting from the same elastic modeling and following basically the same procedure, we can also calculate the SC. The only difference is that the Alexander–Pincus formula (24) is replaced with Eq. (32). In polar coordinates, the calculation of SC reduces to $I_{0,2}$ and $I_{4,2}$ in Appendix B; the result reads
$\frac{\chi_{\gamma}}{S/(\pi\xi_{*}^{2})}=-\left(\vartheta^{2}e^{-\vartheta^{2}}+\frac{\vartheta^{2}e^{-\vartheta^{2}/{\mu_{\text{r}}}}}{{\mu_{\text{r}}}^{2}}\right)\\ {}+\left[Q_{\gamma}(\vartheta^{2})-\dfrac{Q_{\gamma}(\vartheta^{2}/{\mu_{\text{r}}})}{{\mu_{\text{r}}}}\right]\cos{4\varphi_{*}},$ (63)
where we have defined $Q_{\gamma}(w)=(6w^{-1}+4+w)e^{-w}$. This analytical expression is plotted in Fig. 2 as a 2D color map, and delineated in Fig. 5 as a function of $\tilde{{d}}\;({}=2\sqrt{D^{\text{c}}t_{\Delta}}\,\vartheta)$ for $\varphi_{*}=0$ and $\varphi_{*}=\pi/4$.
Obviously, Eq. (63) includes $\cos{4\varphi_{*}}$ in the form of Eq. (36). This $\varphi_{*}$-dependence leads to the well-known cross-shaped “flower” pattern shown in Fig. 2.
The $\xi_{*}$-dependence of Eq. (63) is less obvious. Recalling Eq. (36) to denote the two angular modes with $\chi_{\gamma}^{(0)}$ and $\chi_{\gamma}^{(4)}$, we expand them for small $\vartheta$ to find
$\chi_{\gamma}^{(0)}$ $=\frac{S}{\pi\xi_{*}^{2}}\left[-\frac{1+{\mu_{\text{r}}}^{2}}{{\mu_{\text{r}}}^{2}}\vartheta^{2}+\mathcal{O}(\vartheta^{4})\right]=\mathcal{O}\left(({D^{\text{c}}_{*}}t)^{-1}\right),$ (64)
$\chi_{\gamma}^{(4)}$ $=\frac{S}{\pi\xi_{*}^{2}}\left[\frac{2(1-{\mu_{\text{r}}})}{{\mu_{\text{r}}}}+\mathcal{O}(\vartheta^{6})\right]\simeq\frac{2(1-{\mu_{\text{r}}})S}{\pi{\mu_{\text{r}}}\xi_{*}^{2}};$ (65)
here it should be noted that, since $\tilde{{d}}>\ell_{0}$, the smallness of $\vartheta$ implies largeness of ${D^{\text{c}}_{*}}t$.
It is easy to verify that the above expressions for DCs and SCs are consistently interrelated by Eqs. (45c). In particular, Eq. (61) implies that $f_{+}$ is asymptotically a linear combination of $\{{\ln\xi_{*}},1\}$, which belongs to the null space of $\hat{L}_{+}$ and therefore most of the information of $f_{+}$ is lost from $\chi_{\gamma}^{(0)}$. Contrastively, the information of $f_{-}$ in Eq. (62) is transmitted by $\hat{L}_{-}$ to $\chi_{\gamma}^{(4)}$ properly, in the sense that both $f_{-}$ in Eq. (62) and $\chi_{\gamma}^{(4)}$ in Eq. (65) contain the same factor $(1-{\mu_{\text{r}}})/{\mu_{\text{r}}}$.
VI.2 Logarithmic regime as a reflection of cage effect
We have seen that the logarithmic behavior of the DC tensor can be derived from the elastic modeling. Here the elastic modeling means assumption of approximate dynamics leading to Eqs. (51) for $C_{a}^{\Delta}$ (with $a\in\{\mathrm{d},\mathrm{r}\}$) and allowing derivation of Eq. (61), which implies that the $\xi_{*}$-dependence of $f_{+}$ is asymptotically logarithmic for distances shorter than $2\sqrt{{D^{\text{c}}_{*}}t}$.
Now we will extend this result to a wider class of modeling, having various kinds of viscoelastic or elastoplastic models for ${\psi}_{\mathrm{d}}$ and ${\psi}_{\mathrm{r}}$ in view. Instead of analyzing some specific model in detail, we assume only that $C_{a}^{\Delta}$ derived from the modeling satisfies certain conditions specified below. On this assumption, we will show $f_{+}$ to behave asymptotically as
$\frac{f_{+}}{\sigma^{2}}=-{C_{+}}\ln\frac{\xi_{*}}{\lambda}$ (66)
for $\xi_{*}\gg\lambda$, where $C_{+}$ and $\lambda$ are constant with regard to $\xi_{*}$ but possibly dependent on $t_{\Delta}$. The result of the elastic modeling in Eq. (61) is a special case of Eq. (66).
To proceed, we need some minimal prescription for $C_{a}^{\Delta}=C_{a}^{\Delta}(k,t_{\Delta})$. As is evident from its definition in Eq. (23), the “initial value” of $C_{a}^{\Delta}$ for $t_{\Delta}=0$ is zero, and it is natural to expect that $C_{a}^{\Delta}(k,t_{\Delta})$ is a growing function of $t_{\Delta}$. Taking this $t_{\Delta}$-dependence into account, here we consider the following two types of the $k$-dependence:
  1. 1.
    For shorter wavelengths, $C_{a}^{\Delta}(k,t_{\Delta})$ grows with lapse of $t_{\Delta}$ and reaches a constant saturation value independent of $k$. On the other hand, $C_{a}^{\Delta}(k,t_{\Delta})$ vanishes in the limit of long waves ($k\to 0$).
  2. 2.
    Behaving in proportion to $k^{2}$ over the entire range of wavenumbers, $C_{a}^{\Delta}(k,t_{\Delta})$ diverges for shorter waves and for $t_{\Delta}\to\infty$.
These are two typical behaviors of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$ that we have encountered in Ref. [20]. The first type leads to Eq. (66), while the second type does not. Note that the classification is not comprehensive, but discussion on other types of behavior is out of the scope of the present work.
VI.2.1 First type: caged behavior with a saturation value
The logarithmic behavior in Eq. (66) is shown to result from the first type mentioned above, characterized by the emergence of a constant saturation value. From the viewpoint of particle dynamics, the saturation of $C_{a}^{\Delta}(k,t_{\Delta})$ is understood as reflection of the cage effect on the rotational or dilatational modes of mesoscopic deformation. Note that $\tau_{\alpha}$, usually regarded as representing the lifetime of the cage, is measured with $F_{\text{S}}(k,t_{\Delta})$ which is a single-particle quantity; this means that $\tau_{\alpha}$ measures the collapse of the cage on the lengthscale of $\sigma$, but does not eliminate the possibility that, for other lengthscales, the cage effect may last for times longer than $\tau_{\alpha}$.
The assumptions of the saturation stated above can be reformulated more precisely by postulating the existence of a (nondimensionalized) lengthscale, $\lambda=\lambda(t_{\Delta})$, such that
$C_{a}^{\Delta}(k,t_{\Delta})\sim\begin{cases}D_{*}k^{2}t_{\Delta}&{(0<k\ll 1/\lambda)}\\ {{S_{a}^{\sharp}}}&{({1/\lambda}\ll k<k_{\text{max}})}\end{cases}$ (67)
where ${S_{a}^{\sharp}}$ is the saturation value (independent of $t_{\Delta}$ and $k$), and $k_{\text{max}}$ is the cutoff wavenumber corresponding to the lengthscale ${}\sim{\ell_{0}}$. For the sake of simplicity, however, we treat $k_{\text{max}}$ as if it is infinitely large. It is also assumed that $0\leq{C_{a}^{\Delta}(k,t_{\Delta})}\leq{S_{a}^{\sharp}}$ over the entire range of $k$. We note that Eqs. (51), given by the elastic modeling, satisfy the assumptions in Eq. (67), with $\lambda\sim 2\sqrt{D^{\text{c}}_{*}t_{\Delta}}$, ${{S_{\mathrm{d}}^{\sharp}}}=S$ and ${{S_{\mathrm{r}}^{\sharp}}}=S/{\mu_{\text{r}}}$, and therefore the behavior of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$ given by the elastic modeling belongs to this type.
Unlike Eqs. (51), $C_{a}^{\Delta}(k,t_{\Delta})$ in general is not necessarily reducible to the form tractable with the integrals given in Appendix B. Therefore we must go back to the Alexander–Pincus formula, given as Eqs. (26) in polar coordinates. To denote the integrals contained in Eqs. (26) conveniently, we define
$I_{a}^{(m)}(\xi_{*})=\int_{0}^{\infty}C_{a}^{\Delta}(k,t_{\Delta})J_{m}(k\xi_{*})\frac{\mathrm{d}{k}}{k}$ (68)
where $a\in\{\mathrm{d},\mathrm{r}\}$ and $m$ is an integer; for evaluation of Eqs. (26) we need only $I_{a}^{(0)}$ and $I_{a}^{(2)}$.
To demonstrate the presence of the logarithmic regime, let us evaluate the integral in Eq. (68) for the shorter-scale range of $\xi_{*}$ specified as $k_{\text{max}}^{-1}\ll\xi_{*}\ll\lambda$. In consideration of the factor $1/k$ in the integrand, we split the integral at the wavenumber $1/\lambda$, as
$I_{a}^{(m)}(\xi_{*})=\int_{0}^{1/\lambda}C_{a}^{\Delta}(k,t_{\Delta})J_{m}(k\xi_{*})\frac{\mathrm{d}{k}}{k}\\ {}+\int_{1/\lambda}^{\infty}C_{a}^{\Delta}(k,t_{\Delta})J_{m}(k\xi_{*})\frac{\mathrm{d}{k}}{k}.$ (69)
For $m\geq 1$, it is easy to show that the contribution from the vicinity of $k=0$ is negligible, because $J_{m}(k\xi_{*})$ is small in proportion to $(k\xi_{*})^{m}$ for small $k$. The second integral can be evaluated by changing the variable of integration from $k$ to $q=k\xi_{*}$ and noticing that the lower bound of $q$ equals $\xi_{*}/\lambda\ll 1$. In particular, for $m=2$ we have
$I_{a}^{(2)}(\xi_{*})\simeq{S_{a}^{\sharp}}\int_{1/\lambda}^{\infty}J_{2}(k\xi_{*})\frac{\mathrm{d}{k}}{k}={\frac{1}{2}}{S_{a}^{\sharp}}$ (70)
with the terms of $\mathcal{O}\left({(\xi_{*}/\lambda)^{2}}\right)$ discarded.
The case of $m=0$ requires a more careful treatment. The first integral on the right-hand side of Eq. (69) does not vanish; rather, it is shown to remain finite for small $\xi_{*}$. Taking the $k$-dependence of $C_{a}^{\Delta}(k,t_{\Delta})$ prescribed in Eq. (67) into account, at the lowest order of the longwave approximation, we can evaluate the integral as
$\int_{0}^{1/\lambda}C_{a}^{\Delta}(k,t_{\Delta})J_{0}(k\xi_{*})\frac{\mathrm{d}{k}}{k}\simeq\int_{0}^{1/\lambda}\frac{C_{a}^{\Delta}(k,t_{\Delta})}{k}\mathrm{d}{k}\\ {}\simeq{D_{*}t_{\Delta}}\int_{0}^{1/\lambda}k\,\mathrm{d}{k}=\frac{D_{*}t_{\Delta}}{2\left\{\lambda(t_{\Delta})\right\}^{2}}$ (71)
to find it finite. Consideration of terms in higher order of $k$ in the longwave approximation does not change the conclusion that the integral gives a finite value.
To evaluate the second integral, we make use of $(\ln{k\lambda})^{\prime}=1/k$, with the prime denoting $\partial/\partial{k}$. Upon integration by parts, we have
$\int_{1/\lambda}^{\infty}C_{a}^{\Delta}(k,t_{\Delta})J_{0}(k\xi_{*})\frac{\mathrm{d}{k}}{k}$
$=\int_{1/\lambda}^{\infty}C_{a}^{\Delta}(k,t_{\Delta})J_{0}(k\xi_{*})(\ln{k\lambda})^{\prime}\mathrm{d}{k}$
$\simeq{-{S_{a}^{\sharp}}}\int_{1/\lambda}^{\infty}[J_{0}(k\xi_{*})]^{\prime}(\ln{k\lambda})\mathrm{d}{k}.$ (72)
Subsequently, changing the variable from $k$ to $q=k\xi_{*}$, we find
$[J_{0}(k\xi_{*})]^{\prime}\mathrm{d}{k}=\frac{\mathrm{d}[J_{0}(q)]}{\mathrm{d}q}\mathrm{d}{q}=-J_{1}(q)\mathrm{d}{q},$
$\ln{k\lambda}=\ln\frac{q\lambda}{\xi_{*}}=\ln{q}-\ln{\frac{\xi_{*}}{\lambda}}$
so that
$\int_{1/\lambda}^{\infty}[J_{0}(k\xi_{*})]^{\prime}(\ln{k\lambda})\mathrm{d}{k}$ $=-\int_{\xi_{*}/\lambda}^{\infty}\left(\ln{q}-\ln{\frac{\xi_{*}}{\lambda}}\right)J_{1}(q)\mathrm{d}{q}$
$\simeq\int_{0}^{\infty}\left(\ln{\frac{\xi_{*}}{\lambda}}-\ln{q}\right)J_{1}(q)\mathrm{d}{q}$ (73)
for $\xi_{*}\ll\lambda$. Evaluating the integral and combining the result with Eq. (71), finally we obtain
$I_{a}^{(0)}(\xi_{*})\simeq-{S_{a}^{\sharp}}\ln\frac{\xi_{*}}{\lambda}+\mathcal{O}(1).$ (74)
Note that the term of $\mathcal{O}(1)$ in Eq. (74) includes the contribution of Eq. (71) which is constant with regard to $\xi_{*}$ but may or may not depend on $t_{\Delta}$. Without loss of generality, the term of $\mathcal{O}(1)$ can be set equal to zero, because Eq. (67) leaves room for redefining $\lambda$ to absorb it.
Now suppose that a certain viscoelastic modeling under consideration gives caged dynamics, both for ${C}_{\mathrm{d}}^{\Delta}$ and for ${C}_{\mathrm{r}}^{\Delta}$, so that Eqs. (70) and (74) are valid for both modes. Using Eqs. (70) and (74), we can evaluate the DC tensor given by Eqs. (26), from which we can extract $f_{\pm}$ by Eqs. (40) and (41). The result reads
$f_{+}$ $\simeq-\frac{\ell_{0}^{2}}{\pi}\left({S_{\mathrm{d}}^{\sharp}}+{S_{\mathrm{r}}^{\sharp}}\right)\ln\frac{\xi_{*}}{\lambda},$ (75a)
$f_{-}$ $\simeq\frac{\ell_{0}^{2}}{2\pi}\left(-{S_{\mathrm{d}}^{\sharp}}+{S_{\mathrm{r}}^{\sharp}}\right).$ (75b)
Thus $f_{+}$ is found to behave logarithmically as in Eq. (66), with $C_{+}=({\ell_{0}^{2}}/(\pi\sigma^{2}))({S_{\mathrm{d}}^{\sharp}}+{S_{\mathrm{r}}^{\sharp}})$.
As was noted earlier in connection with Eq. (67), the elastic modeling in the previous subsection is a special case of caged dynamics satisfying Eq. (67). It is easy to confirm that setting ${{S_{\mathrm{d}}^{\sharp}}}=S$ and ${{S_{\mathrm{r}}^{\sharp}}}=S/{\mu_{\text{r}}}$ in Eqs. (75a) and (75b), with $\lambda\sim 2\sqrt{{D^{\text{c}}_{*}}t_{\Delta}}$, reproduces Eqs. (61) and (62).
VI.2.2 Second type: uncaged behavior for shorter waves
As a contrast to the caged behavior characterized by Eq. (67), here we consider another type of $k$-dependence, in which $C_{a}^{\Delta}(k,t_{\Delta})$ diverges without being saturated by cage effect. More specifically, we assume
$C_{a}^{\Delta}(k,t_{\Delta})\simeq D_{*}k^{2}t_{\Delta}$ (76)
for the entire range of $k$, with the case of $a=\mathrm{r}$ in mind. This type of behavior is seen in ${C}_{\mathrm{r}}^{\Delta}$ when the fluctuating dynamics of ${\psi}_{\mathrm{r}}$ are equivalent to those of free Brownian motion without restoring force, as is the case in the linear equation of ${\psi}_{\mathrm{r}}$ studied in Subsec. IV-C of Ref. [20].
The integrals needed for evaluation of the DC tensor are still given by Eq. (68) with $m=2$ and $m=4$. For $C_{a}^{\Delta}$ now we use Eq. (76), which implies oscillatory divergence of the integrand for $k\to\infty$. Handling the divergence with the standard technique of convergence factor we find that $I_{a}^{(0)}(\xi_{*})$ vanishes for $\xi_{*}>0$ (the integral reduces to the delta function of $\xi_{*}$). For $m=2$, using the convergence factor $e^{-\epsilon{k}}$ and taking the limit of $\epsilon\to+0$, we obtain
$\int_{0}^{\infty}{C}_{\mathrm{r}}^{\Delta}(k,t-s)J_{2}(k\xi_{*})e^{-\epsilon{k}}\frac{\mathrm{d}{k}}{k}\\ {}=D_{*}\,(t-s)\int_{0}^{\infty}J_{2}(k\xi_{*})e^{-\epsilon{k}}k\mathrm{d}{k}\to\frac{2D_{*}\,(t-s)}{\xi_{*}^{2}},$ (77)
which diverges for $\xi_{*}\to 0$.
The results obtained above can be applied to the case of linearized dynamics studied in Subsec. IV-C of Ref. [20]. In this case, the rotational mode is uncaged so that ${C}_{\mathrm{r}}^{\Delta}$ behaves according to Eq. (76), while ${C}_{\mathrm{d}}^{\Delta}$ is subject to saturation as in Eq. (67). Taking $I_{\mathrm{r}}^{(0)}(\xi_{*})=0$ into account, we find
$f_{+}$ $\simeq-\frac{\ell_{0}^{2}}{\pi}{S_{\mathrm{d}}^{\sharp}}\ln\frac{\xi_{*}}{\lambda},$ (78a)
$f_{-}$ $\simeq\frac{2{\ell_{0}^{2}}D_{*}t_{\Delta}}{\pi{\xi_{*}^{2}}}$ (78b)
in this case. The result is consistent with Eq. (4.37) of Ref. [20].
VI.3 Quantitative data analysis of the logarithmic regime
With Eqs. (75) and (78) relating $f_{\pm}$ to the two types of behavior of $C_{a}^{\Delta}$, now let us analyze the numerical data of $f_{\pm}$. The analysis may allow us to extract information of the assumed dynamics, such as the elastic moduli.
(a)Refer to caption(b)Refer to caption
Figure 12: $f_{-}$ nondimensionalized with $\sigma^{2}$ and plotted against $\xi_{*}=\tilde{{d}}/\ell_{0}$. (a) Plots in the case of $\phi_{\text{area}}=0.5$ for three different values of the time interval: $t=40$, $80$ and $160$. (b) Plots in the cases of $\phi_{\text{area}}=0.6$ and $\phi_{\text{area}}=0.7$, for $t=80$ (closed symbols) and $t=160$ (open symbols).
The main question here is whether ${C}_{\mathrm{r}}^{\Delta}$ is caged or uncaged. This is answered by checking which of Eq. (75b) and Eq. (78b) is closer to the actual behavior of $f_{-}$. From Fig. 12 where $f_{-}$ is plotted against $\xi_{*}=\tilde{{d}}/\ell_{0}$ , we can judge that ${C}_{\mathrm{r}}^{\Delta}$ is caged for all the three values of $\phi_{\text{area}}$ studied here. As is shown in Fig. 12(a), the values of $f_{-}$ in the case of $\phi_{\text{area}}=0.50$ is nearly constant (i.e. independent of $\xi_{*}$ and $t_{\Delta}$), except for deviation at very small distance. We see from Fig. 12(b) that $f_{-}$ for $\phi_{\text{area}}=0.60$ is also nearly constant. The constancy of $f_{-}$ seems to be consistent with Eq. (75b). The case of $\phi_{\text{area}}=0.70$ is difficult to interpret, as $f_{-}$ is nor constant nor proportional to $\xi_{*}^{-2}$, but at least we can eliminate the divergent behavior in Eq. (78b) predicted for the case of uncaged ${C}_{\mathrm{r}}^{\Delta}$. Thus we can regard Eqs. (75) as a reasonable approximation of $f_{\pm}$, which provides supporting evidence that the cage effect may last longer than $\tau_{\alpha}$ at some lengthscales greater than $\ell_{0}$.
As a corollary of the above result, we find that $f_{+}$ dominates over $f_{-}$ for $\xi_{*}\ll\lambda$. On the basis of this dominance, we can estimate the magnitude of the DC at the typical interparticulate distance $\ell_{0}$, by extrapolating the asymptotic behavior of $f_{+}$ in Eq. (75a) to $\xi_{*}=1$. This is approximately the value of MSD that the particle would have if it were eternally confined in the cage of the neighboring particles. Denoting it with $\left\langle{\mathbf{R}^{2}}\right\rangle_{\text{caged}}$, for $\lambda=\lambda(t)\gg 1$ we have
$\left\langle{\mathbf{R}^{2}}\right\rangle_{\text{caged}}\approx\left.{f_{+}}\right|_{{\xi_{*}}=1}\simeq\frac{\ell_{0}^{2}}{\pi}\left({S_{\mathrm{d}}^{\sharp}}+{S_{\mathrm{r}}^{\sharp}}\right)\ln{\lambda(t)},$ (79)
which gives the time-dependent version of the Mermin–Wagner fluctuation,
$\left\langle{\mathbf{R}^{2}}\right\rangle_{\text{caged}}\propto\frac{\ell_{0}^{2}}{2\pi}\left({S_{\mathrm{d}}^{\sharp}}+{S_{\mathrm{r}}^{\sharp}}\right)\ln({D_{*}}t_{\Delta}),$ (80)
if $\lambda(t)$ behaves as $\sqrt{{D_{*}}t_{\Delta}\mathstrut}$ and is greater than unity but smaller than the system size.
Figure 13: ${C_{+}=-\mathrm{d}({f_{+}/\sigma^{2}})/\mathrm{d}(\ln\xi_{*})}$ for $\phi_{\text{area}}=0.5$, $0.6$ and $0.7$, plotted against $t$.
Figure 14: Length scale $\lambda$ evaluated by fitting Eq. (66) to the numerical data of $f_{+}=X_{\parallel}+X_{\perp}$. The solid line indicates $\lambda\propto t_{\Delta}^{1/2}$.
Figure 15: The values of ${f_{-}\simeq{\frac{1}{2}}{\sigma^{2}}{C_{-}}}$ at $\xi_{*}=1$ for $\phi_{\text{area}}=0.5$, $0.6$ and $0.7$, plotted against $t$, from which $C_{-}$ is estimated.
Next, we discuss how to extract quantitative information from the data of $f_{+}$ by means of Eq. (75a). More specifically, we aim to evaluate ${S_{\mathrm{r}}^{\sharp}}$, which is expected to carry information of the shear modulus. Noticing that Eq. (75a) is in the form of Eq. (66), we fit this form to the data of $f_{+}$ as a function of $\xi_{*}$, within the range in which the logarithmic behavior is observed, to obtain $C_{+}=-\mathrm{d}(f_{+}/\sigma^{2})/\mathrm{d}(\ln\xi_{*})$ and $\lambda$ as fitting parameters. In other words, we regard Eq. (66) as the definition of $C_{+}$ and $\lambda$ through fitting. The values of $C_{+}$ and $\lambda$ thus obtained for various time intervals, $t_{\Delta}$, are plotted in Fig. 13 and Fig. 14, respectively.
Let us examine the values of $C_{+}$ in Fig. 13. For $\phi_{\text{area}}=0.5$ and $\phi_{\text{area}}=0.6$, $C_{+}$ is nearly independent of $t_{\Delta}$. In the case of $\phi_{\text{area}}=0.7$, $C_{+}$ grows as $t_{\Delta}$ increases; the limitation due to the finite system size makes it difficult to determine whether $C_{+}$ converges to some limiting value or grows unlimitedly. If it actually converges, we may say that the limiting value of $C_{+}$ increases as $\phi_{\text{area}}$ is increased.
From the logarithmic plot in Fig. 14, we can read that the length scale $\lambda=\lambda(t_{\Delta})$ grows nearly in proportion to $\sqrt{t_{\Delta}}$. This diffusive behavior of $\lambda(t_{\Delta})$ is consistent with the prediction of the elastic model that the DCs are expressed in terms of the similarity variable $\vartheta$ in Eq. (50). It also supports one of the assumptions underlying Eq. (80) about $\left\langle{\mathbf{R}^{2}}\right\rangle_{\text{caged}}$. In regard to the $\phi_{\text{area}}$-dependence of $\lambda$, we find that $f_{+}$ becomes less diffusive as $\phi_{\text{area}}$ is increased.
Lastly, we ask whether quantitative information extracted from the data is consistent with the elastic modeling. Such information must be contained, at least partially, in the values of $C_{+}$ obtained from Fig. 13. Since $C_{+}$ equals $(\ell_{0}^{2}/(\pi\sigma^{2}))({S_{\mathrm{d}}^{\sharp}}+{S_{\mathrm{r}}^{\sharp}})$ according to Eq. (75a), it should be possible to obtain therefrom the value of ${S_{\mathrm{r}}^{\sharp}}$, which is related to the shear modulus, if ${S_{\mathrm{d}}^{\sharp}}$ is known somehow. There are at least two possible ways to evaluate ${S_{\mathrm{d}}^{\sharp}}$: we may substitute for it the static structure factor of the density, $S$, or we may use Eq. (75b) to estimate $-{S_{\mathrm{d}}^{\sharp}}+{S_{\mathrm{r}}^{\sharp}}$ on the basis of $f_{-}$.
Let us discuss the first choice (i.e. evaluation from $S$ and $C_{+}$). We can estimate
${S_{\mathrm{r}}^{\sharp}}\approx\frac{\pi\sigma^{2}}{\ell_{0}^{2}}{C_{+}}-S=4\phi_{\text{area}}{C_{+}}-S$ (81)
from the value of $S$ in Table 1 and $C_{+}$ read from Fig. 13. For $\phi_{\text{area}}=0.50$, we find ${S_{\mathrm{r}}^{\sharp}}$ to be vanishing small; although literal interpretation of Eq. (81) gives ${S_{\mathrm{r}}^{\sharp}}=-0.02$, we must remember that ${S_{\mathrm{r}}^{\sharp}}$ cannot be negative. In the case of $\phi_{\text{area}}=0.60$ we have ${S_{\mathrm{r}}^{\sharp}}\approx 0.6$ according to Eq. (81), and ${S_{\mathrm{r}}^{\sharp}}\approx 2.8$ if we take $C_{+}\approx 1$ for $\phi_{\text{area}}=0.70$.
The second choice turns out to be perplexing. In view of the approximate constancy of $f_{-}$ within a certain range of $\xi_{*}$ shown in Fig. 12, we write the constant as $f_{-}\simeq({\sigma^{2}}/2)C_{-}$ in parallel with $C_{+}$. The values of $C_{-}$, shown in Fig. 15, are nearly independent of $t_{\Delta}$ (except for the decay after $t=70$ in the case of $\phi_{\text{area}}=0.5$). In terms of $C_{-}$ thus defined, Eq. (75b) reads $C_{-}\simeq({\ell_{0}^{2}}/(\pi\sigma^{2}))(-{S_{\mathrm{d}}^{\sharp}}+{S_{\mathrm{r}}^{\sharp}})$ , so that Eqs. (75) yield
${S_{\mathrm{d}}^{\sharp}}$ $\approx\frac{\pi\sigma^{2}}{2\ell_{0}^{2}}\left({C_{+}}-{C_{-}}\right)=2\phi_{\text{area}}\left({C_{+}}-{C_{-}}\right),$ (82a)
${S_{\mathrm{r}}^{\sharp}}$ $\approx\frac{\pi\sigma^{2}}{2\ell_{0}^{2}}\left({C_{+}}+{C_{-}}\right)=2\phi_{\text{area}}\left({C_{+}}+{C_{-}}\right).$ (82b)
Since both ${S_{\mathrm{d}}^{\sharp}}$ and ${S_{\mathrm{r}}^{\sharp}}$ must be positive, we should have $0<C_{-}<C_{+}$, but the values of $C_{-}$ read from Fig. 15 ($C_{-}\simeq 2f_{-}/\sigma^{2}\approx 0.4$ both for $\phi_{\text{area}}=0.50$ and $\phi_{\text{area}}=0.60$) contradict this inequality, as the value is greater than $C_{+}$ in Fig. 13 ($C_{+}<0.1$ for $\phi_{\text{area}}=0.5$ and $C_{+}\approx 0.3$ for $\phi_{\text{area}}=0.6$). In the case of $\phi_{\text{area}}=0.70$, the “second choice” estimation from $C_{\pm}$ by way of Eq. (82b) gives ${S_{\mathrm{r}}^{\sharp}}\approx 2.0$ (if we take $C_{+}\approx 1$ and $C_{-}\approx 0.4$), which is somewhat smaller than ${S_{\mathrm{r}}^{\sharp}}\approx 2.6$ obtained from $S$ and $C_{+}$ (the first choice).
In spite of the quantitative inconsistency, however, the two ways of estimating ${S_{\mathrm{r}}^{\sharp}}$ as a function of $\phi_{\text{area}}$ are in agreement about qualitative tendency. As $\phi_{\text{area}}$ is increased from $0.6$ to $0.7$, the value of ${S_{\mathrm{r}}^{\sharp}}$ estimated by Eq. (81) increases from about $0.6$ to $2.6$, and also the estimation by Eq. (82b) shows an increase from $0.9$ to $2.0$. This tendency suggests the need of something more sophisticated than the simple elastic modeling, because, according to Eq. (51b) derived from Eq. (52b), ${S_{\mathrm{r}}^{\sharp}}$ is inversely proportional to the shear modulus and therefore supposed to be a decreasing function of $\phi_{\text{area}}$. The present analysis, revealing that ${S_{\mathrm{r}}^{\sharp}}$ is an increasing function of $\phi_{\text{area}}$ for $\phi_{\text{area}}\leq 0.7$ and $t_{\Delta}\gg\tau_{\alpha}$, provides useful information for searching suitable models about the dynamics of ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$.
VII Concluding remarks
We have investigated three questions about space–time correlations in a model colloidal liquid. First, we asked whether the displacements have nontrivial correlations even in liquids that are only slightly glassy. The answer is affirmative: this is demonstrated by computing $\mathsf{X}$, the DC tensor defined in Eq. (5), and $\chi_{\gamma}$, the SC defined by Eq. (12), from simulation data of a two-dimensional model colloidal liquid with $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$. The time dependence of these correlations is not linear with regard to the time interval $t_{\Delta}$, as would be expected if the detected correlations were explained as accumulation of many events, but rather described in terms of a similarity variable $\vartheta$ in Eq. (50), indicating the presence of the diffusive correlation length.
Secondly, we asked whether DC and SC are equivalent. The answer is negative: to answer this question, we have derived Eqs. (45c) as a relation between the two kinds of correlations, by treating the displacement field with the label variable formulation. The computed values of DC and SC are then shown to be consistent with Eqs. (45c).
The relation in Eqs. (45c) takes the form of a linear mapping from the DC to the SC, which is expressed by means of two linear operators, $\hat{L}_{\pm}$, of Euler–Cauchy type. This mapping is non-invertible: we can discover $\chi_{\gamma}$ from the components of the DC, $f_{\pm}=X_{\parallel}\pm X_{\perp}$, but cannot recover $f_{\pm}$ uniquely from $\chi_{\gamma}$. In particular, if $f_{+}$ falls into the null space of $\hat{L}_{+}$, the information of $f_{+}$ is lost due to the mapping. In this sense, the DC and the SC are not equivalent.
Noticing that the null space of $\hat{L}_{+}$ is spanned by $\{1,\log\xi_{*}\}$, we have found from the simulation data that $f_{+}$ indeed behaves like $\ln\xi_{*}$ for shorter distances. The information of this logarithmic behavior in the DC is therefore lost from the SC. As a result, the SC has weaker dependence on $\phi_{\text{area}}$ than the DC.
Thus we have evidenced that the DC is more informative than the SC. The extra information can be obtained from the numerical data by fitting Eq. (66) to the asymptotic logarithmic behavior of $f_{+}$.
Interpretation of this information is our third question. It is manifestation of the cage effect at spatiotemporal scales other than usually noticed. The logarithmic behavior of the DC can be explained by assuming a kind of “caged” dynamics for $C_{a}^{\Delta}$, such that the cage effect survives for times longer than $\tau_{\alpha}$ at lengthscales greater than $\ell_{0}$, so that temporal growth of $C_{a}^{\Delta}$ is saturated as is prescribed in Eq. (67). A simple version of such dynamics is exemplified by the elastic modeling in Eq. (52). The dynamics assumed behind Eq. (67), including the elastic modeling, can account for the logarithmic behavior of $f_{+}$ and the non-divergent behavior of $f_{-}$ at the shorter lengthscales. The elastic modeling, however, turns out to be too simple to reproduce the $\phi_{\text{area}}$-dependence of the coefficient $C_{+}$ of the logarithmic behavior in Eq. (66): while the numerical data suggests that $C_{+}$ is an increasing function of $\phi_{\text{area}}$, the elastic modeling makes the opposite prediction, as was noted toward the end of Subsec. VI.3.
As the logarithmic behavior of $f_{+}$ comes from $\hat{L}_{+}$, which is essentially the 2D Laplacian, its counterpart in other dimensionalities (1D and 3D) deserves some remarks. The DC in the 1D setup [40, 26] is known to behave as
$\left\langle{{R_{i}}{R_{j}}}\right\rangle=2S\ell_{0}\sqrt{\frac{D^{\text{c}}_{*}{t}}{\pi}}\left(e^{-\vartheta^{2}}-{\sqrt{\pi}}\left|{\vartheta}\right|\operatorname{erfc}\left|{\vartheta}\right|\right),$ (83)
whose shorter-distance asymptotic form, $\left\langle{{R_{i}}{R_{j}}}\right\rangle\propto 1-{\sqrt{\pi}}\left|{\vartheta}\right|$, is essentially a fundamental solution to the 1D Laplacian. In the 3D case, we expect that the DC in some conditions may behave like the fundamental solution to the 3D Laplacian, namely $1/\vartheta$. Numerical exploration of such behavior would be an interesting direction of investigation.
Let us now remark on what could be done in the near future within the 2D setup. Since $C_{+}$ seems to carry information on some long-lived aspect of the cage effect at lengthscales greater than $\ell_{0}$, it will be desirable to develop more sophisticated modeling for the dynamics of $C_{a}^{\Delta}$ at such scales. Prediction of $C_{+}$ as a function of $\phi_{\text{area}}$ will serve as a touchstone of such modeling. Viscoelastic modelings of mesoscopic deformations by combination of Maxwell-like phenomenology and the mode-coupling theory (MCT) would be a hopeful direction. While the recent work by Maier et al. in this direction is based on Mori–Zwanzig projection formalism of momentum balance [41], we have been developing a different version of MCT for ${C}_{\mathrm{d}}^{\Delta}$ and ${C}_{\mathrm{r}}^{\Delta}$ on the basis of the (overdamped) Dean–Kawasaki equation [42, 20]. Details on this version of MCT will be reported elsewhere.
As another possible direction to extend the present work, we may mention cases of higher $\phi_{\text{area}}$, which makes the liquid more glassy and the relaxation time long enough to allow comparison between the plateau regime, $\tau_{0}\ll t_{\Delta}\ll\tau_{\alpha}$, and the “post-$\tau_{\alpha}$” regime, $t_{\Delta}\gg\tau_{\alpha}$. The elastic model is expected to be valid in the plateau regime and therefore we may be able to extract information of the elastic moduli from the logarithmic behavior of the DC. The “post-$\tau_{\alpha}$” regime with long $\tau_{\alpha}$ and high $\phi_{\text{area}}$ is where Doliwa and Heuer [11] found the significance of the directional aspect in studying space–time correlations in glassy liquids. Quantitative analysis of $f_{\pm}$ in such cases will provide deep insights into caged dynamics, hopefully even more valuable than the impressive pictures of the DC (Fig. 1) and the SC (Fig. 2).
Acknowledgements.
We express our cordial gratitude to Grzegorz Szamel, Erika Eiser, Peter Schall, Hisao Hayakawa, Norihiro Oyama, Kunimasa Miyazaki, Atsushi Ikeda, Takenobu Nakamura, Hajime Yoshino, Kang Kim, Takahiro Hatano, and So Kitsunezaki for fruitful discussions, insightful comments and valuable suggestions. The first author (Ooshida) also thank Eni Kume and Shiladitya Sengupta for informative discussions on the occasion of recent international conferences held in the difficult time of the Covid19 pandemic. In addition, the assistance of Susumu Goto as a coauthor of Ref. [20] is gratefully acknowledged, especially in preparation of Fig. 1. This work was supported by Grants-in-Aid for Scientific Research (Kakenhi) JP-15K05213, JP-18K03459 and JP-21K03404, JSPS (Japan).
Appendix A Numerical procedure for calculating DC from particle data
The DC tensor defined by Eq. (5) has essentially only two components, denoted with $X_{\parallel}$ and $X_{\perp}$, as is seen in Eq. (6). They are computed by extending the numerical method in Appendix A of Ref. [26] to the 2D system. Here we describe some details of formulation and computation.
Suppose that we have data of $\{\mathbf{r}_{i}(t)\}_{i=1,2,\ldots,N}$, from which we prepare the displacement vector $\mathbf{R}_{i}=\mathbf{R}_{i}(t,s)$ according to Eq. (1), and the relative position vector
$\mathbf{r}_{ij}(s)=\mathbf{r}_{j}(s)-\mathbf{r}_{i}(s)$
for all pairs $(i,j)$ within some distance (at most half the system size $L$). On the basis of these data, the DC tensor is prescribed in Eq. (5) by means of the conditional average, $\left\langle{\quad}\right\rangle_{\tilde{\mathbf{d}}}$.
The conditional average is conceptually formulated as
$\left\langle{O}\right\rangle_{\tilde{\mathbf{d}}}=\frac{\sum_{i,j}\left\langle{O_{ij}\delta(\mathbf{r}_{ij}(s)-\tilde{\mathbf{d}})}\right\rangle}{\sum_{i,j}\left\langle{\delta(\mathbf{r}_{ij}(s)-\tilde{\mathbf{d}})}\right\rangle}$ (84)
for any physical observable $O_{ij}$ associated with the particle pair $(i,j)$. In the case of the DCs, as the observable $O_{ij}$ we take
$(X_{\parallel})_{ij}$ $=({\mathbf{e}_{ij}}\cdot\mathbf{R}_{i})({\mathbf{e}_{ij}}\cdot\mathbf{R}_{j}),$ (85)
$(X_{\perp})_{ij}$ $=\det(\mathbf{e}_{ij},{\mathbf{R}_{i}})\det(\mathbf{e}_{ij},{\mathbf{R}_{j}}),$ (86)
using orthogonal decomposition of the displacement vectors with regard to
$\mathbf{e}_{ij}=\frac{\mathbf{r}_{ij}(s)}{\left|{\mathbf{r}_{ij}(s)}\right|}.$ (87)
To evaluate $X_{\parallel}(\tilde{{d}},t_{\Delta})=\left\langle{X_{\parallel}}\right\rangle_{\tilde{\mathbf{d}}}$ and $X_{\perp}(\tilde{{d}},t_{\Delta})=\left\langle{X_{\perp}}\right\rangle_{\tilde{\mathbf{d}}}$ numerically, we approximate the delta function in Eq. (84) by a statistical bin with $\Delta{r}$ in width (we use $\Delta{r}=L/200\approx\sigma/3$ for the present calculations with $N=4000$), assuming the statistical isotropy of the system at once. Thus Eq. (84) is discretized as
$\left\langle{O}\right\rangle_{\tilde{\mathbf{d}}}=\frac{\left\langle{\sum_{ij}O_{i,j}\Theta(\tilde{{d}}\leq\left|{\mathbf{r}_{ij}(s)}\right|<\tilde{{d}}+\Delta{r})}\right\rangle}{\left\langle{\sum_{i,j}\Theta(\tilde{{d}}\leq\left|{\mathbf{r}_{ij}(s)}\right|<\tilde{{d}}+\Delta{r})}\right\rangle},$ (88)
where $\Theta$ denotes the indicator function of the statistical bin, such that its value equals unity if and only if $\left|{\mathbf{r}_{ij}(s)}\right|$ satisfies the inequality and otherwise zero. Finally, with the ensemble average that remains in Eq. (88) taken over many runs, we obtain $X_{\parallel}$ and $X_{\perp}$ numerically.
Appendix B Wavenumber integrals involving Bessel function and Gaussian
For analytical calculations, we define
$I_{m,n}(\xi,\tau)=\int_{0}^{\infty}e^{-{\tau}k^{2}}J_{m}(k\xi)\,k^{n-1}\mathrm{d}{k}$ (89)
where $(m,n)$ is a pair of non-negative integers, $\xi$ and $\tau$ are positive real numbers, and $J_{m}$ denotes the Bessel function of order $m$. Evaluation of this integral is required in Subsec. VI.1 for $(m,n)=(0,2)$, $(2,2)$ and $(4,2)$.
Let us begin with $I_{0,2}(\xi,\tau)$. Making use of well-known relations between $J_{0}$ and $J_{1}$ [32], we find
$2\tau I_{0,2}(\xi,\tau)=1-\xi I_{1,1}(\xi,\tau),$ (90)
$\frac{\partial}{\partial\xi}\left[\xi I_{1,1}(\xi,\tau)\right]=\xi I_{0,2}(\xi,\tau),$ (91)
which yields
$2\tau\frac{\partial I_{0,2}(\xi,\tau)}{\partial\xi}=-{\xi}I_{0,2}(\xi,\tau)$ (92)
upon elimination of $I_{1,1}(\xi,\tau)$. This is readily integrated to give
$I_{0,2}(\xi,\tau)=\frac{1}{2\tau}\exp\left({-\frac{\xi^{2}}{4\tau}}\right)$ (93)
with the initial condition, $I_{0,2}(0,\tau)=1/(2\tau)$, taken into account.
Subsequently, in order to evaluate $I_{2,2}(\xi,\tau)$, we notice that the derivative of $J_{1}$ can be expressed in terms of $J_{0}$ and $J_{2}$, which allows us to find
$2\frac{\partial I_{1,1}(\xi,\tau)}{\partial\xi}=I_{0,2}(\xi,\tau)-I_{2,2}(\xi,\tau).$ (94)
Combining Eq. (94) with Eq. (90), after some calculation, we have
$I_{2,2}(\xi,\tau)={\frac{2}{\xi^{2}}}-\frac{1}{2\tau}\left(1+\frac{4\tau}{\xi^{2}}\right)\exp\left(-\frac{\xi^{2}}{4\tau}\right).$ (95)
In an analogous way, we also find
$I_{4,2}(\xi,\tau)=\frac{4}{\xi^{2}}\left(1-\frac{12\tau}{\xi^{2}}\right)\\ {}{}+\frac{1}{2\tau}\left(1+\frac{16\tau}{\xi^{2}}+\frac{96\tau^{2}}{\xi^{4}}\right)\exp\left(-\frac{\xi^{2}}{4\tau}\right),$ (96)
making further usage of the recurrence relations for the Bessel function.

### v2

@@
 The formulations mentioned above are focused on the mobility of particles and not on their direction of motion.
-The importance of the directional aspect of motion was noticed by Doliwa and Heuer [11], who demonstrated that the displacement correlations in the longitudinal direction behave quite differently from those in the transverse direction.
+The importance of the directional aspect of motion was noticed by Doliwa and Heuer [11], who demonstrated with Monte Carlo simulations that the displacement correlations in the longitudinal direction behave quite differently from those in the transverse direction.
+Essentially the same spatial pattern of displacement correlations was also observed independently by Cui et al.
+[12] in experiments with aqueous suspension of silica spheres.
 Later, it was noticed that the transverse displacement correlations contain information about glass elasticity [4, 13, 14].
@@
 Motivated by the above considerations, here we concretize our interest in displacement-based statistical quantities by discussing two-particle displacement correlations (DC) and shear strain correlations (SC) in two-dimensional liquids.
-Our first question is whether and to what extent these correlations are detectable in liquids that are only slightly dense.
+As is exemplified by the aforementioned experiment by Cui et al.
+[12], DC and SC can be measured experimentally.
+Measurement of DC between tracer particles in viscoelastic media has been applied to various soft materials by the name of “two-point micro rheology” [21, 22].
+DCs of directly interacting colloidal particles in a suspension with finite density have also been reported [23].
+Besides, Illing et al.
+[20] conducted video microscopy experiments on SC in a glassy liquid.
+To connect these experimentally (and computationally) measurable statistical quantities with the idea of solid-like liquid theory, now we raise three questions about DC and SC.
+Our first question is whether and to what extent these correlations, namely DC and SC, are detectable in liquids that are only slightly dense.
+To demonstrate the presence of these correlations in liquids is to legitimate the solid-based approach to the liquid dynamics.
 We focus on such density that the lifetime of “bonds” between the nearest neighbors is barely non-zero and not very long.
@@
 The time dependence of these correlations will also be investigated.
+To make better use of DC and SC as experimentally measurable quantities, we ask further questions about their nature.
 The second question concerns the possible equivalence between the DC and the SC, as both of them are calculable from the same data of particle displacements.
@@
 In deriving the Navier–Stokes equation for Newtonian fluids, any confusion between the velocity of a fluid element and the molecular velocity should be severely criticized.
-In dense liquids, however, the cage effect makes it possible—probably as a crude approximation—to regard the velocity of a particle as the velocity of a “fluid element” comprising the particle, with the understanding that fluctuations at the collisional frequency should be filtered out mostly by time-averaging.
+In dense liquids, however, the cage effect makes it possible—probably as a crude approximation—to regard the velocity of a particle as the velocity of a “fluid element” containing the particle, with the understanding that fluctuations at the collisional frequency should be filtered out mostly by time-averaging.
 A simple way to remove the fast fluctuations in the velocity is to focus on the displacement of the particle [28].
@@
 This difference was noted in a pioneering work by Doliwa and Heuer [11], many years before the researchers started to notice that the transverse DC may contain information about shear modulus of glassy systems [14, 13].
+Essentially the same pattern of displacement correlations is reported, seemingly independently, as an experimental result by Cui et al.
+[12], who pointed out the presence of “antidrag” region (where $X_{\perp}<0$ in our notation).
 The directional behavior of DCs can be illustrated pictorially as a pair of vortices [24, 11], as is exemplified in Fig.
@@
 The four-petaled SC pattern is often interpreted as a sign of elasticity, showing the response of an elastic media to localized plastic events.
-Although the observations of the SC pattern in sheared glasses [17, 18] seems to be consistent with this interpretation, it is curious that non-vanishing SC is reported also in quiescent glassy liquids, being visible even for timescales longer than $\tau_{\alpha}$ [19].
+Although the observations of the SC pattern in sheared glasses [18, 19] seems to be consistent with this interpretation, it is curious that non-vanishing SC is reported experimentally also in quiescent glassy liquids, being visible even for timescales longer than $\tau_{\alpha}$ [20].
 Before studying the behavior of SC for $t_{\Delta}\gg\tau_{\alpha}$, here we review how to calculate the SC from the particle simulation data.
@@
 In order to answer this question, let us develop a theoretical framework to derive an analytical relation between the DC and the SC.
+Readers who are more interested in the relation itself than its derivation procedure may skim the theoretical development in Subsec.
+III.1 and, after checking some basic notation in and around Eqs.
+(22)–(25), (36) and (41), jump to Eq.
+(44).
 III.1 Label variable and deformation gradient tensor field
@@
 $\chi_{\gamma}$ $=\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})][\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime})]}\right\rangle$
-$=-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{1^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{y}}(\bm{\xi}^{\prime})}\right\rangle$
-$\qquad{}-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{2^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{x}}(\bm{\xi}^{\prime})}\right\rangle$ (38)
+$=-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{1^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{y}}(\bm{\xi}^{\prime})}\right\rangle-\ell_{0}^{-2}\left\langle{\vphantom{\sqrt{\mathstrut}}\left\{\partial_{2^{\prime}}{[\partial_{1^{\prime}}{R_{y}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})+\partial_{2^{\prime}}{R_{x}}(\bm{\xi}^{\prime}+\bm{\xi}_{*})]}\right\}{R_{x}}(\bm{\xi}^{\prime})}\right\rangle$ (38)
 where $\partial_{\alpha^{\prime}}$ stands for $\partial/\partial{\xi^{\prime}_{\alpha}}$ ($\alpha\in\{1,2\}$). In Eq. (38), use is made of “integration by parts”
@@
 With the relation in Eqs.
-(45c) we can calculate the SC from the DC, but this does not necessarily mean that SC and DC are equivalent.
+(45) we can calculate the SC from the DC, but this does not necessarily mean that SC and DC are equivalent.
 The problem is whether the relation is invertible, making it possible to obtain the DC from the SC.
@@
 The invertibility of the linear operators $\hat{L}_{\pm}$ in Eq.
-(45c) can be discussed by checking their null spaces.
+(45) can be discussed by checking their null spaces.
 The null space of $\hat{L}_{+}$, also known as the kernel of the linear operator [39], is the set of all the solutions $h_{+}$ to the homogeneous equation
@@
 If $f_{+}\sim h_{+}$, it is mapped to zero by $\hat{L}_{+}$ in Eqs.
-(45c) and therefore the information of $f_{+}$ is lost from $\chi_{\gamma}^{(0)}$.
+(45) and therefore the information of $f_{+}$ is lost from $\chi_{\gamma}^{(0)}$.
 This makes the DC and the SC nonequivalent.
@@
 If $f_{-}$ happens to fall into the null space of $\hat{L}_{-}$, the information of $f_{-}$ drops out of $\chi_{\gamma}^{(4)}$.
-Thus we have derived analytical relations involving the SCs, which allows us to discuss the possibility that the DCs and the SCs are not equivalent due to the null space of $\hat{L}$.
+Thus we have derived analytical relations involving the SCs, which allows us to discuss the possibility that the DCs and the SCs are not equivalent due to the null space of $\hat{L}_{\pm}$.
 Now let us proceed to numerical study of these analytical results, beginning with specification of the particle system in the next section.
@@
 $𝟙\left\langle{\mathbf{f}_{i}(t)\otimes\mathbf{f}_{j}(t^{\prime})}\right\rangle=2D\delta_{ij}\delta(t-t^{\prime})\openone.$
-As the interaction potential $V_{jk}$ between the $j$-th and $k$-th particles (separated by the relative position vector $\mathbf{r}_{ij}$), we adopt the harmonic repulsive potential,
-$V_{jk}=\begin{cases}V_{\text{max}}\left(1-\dfrac{\left|{\mathbf{r}_{ij}}\right|}{\sigma}\right)^{2}&(\left|{\mathbf{r}_{ij}}\right|<\sigma)\cr 0&(\text{otherwise})\end{cases}$
+As the interaction potential $V_{jk}$ between the $j$-th and $k$-th particles (separated by the relative position vector $\mathbf{r}_{jk}$), we adopt the harmonic repulsive potential,
+$V_{jk}=\begin{cases}V_{\text{max}}\left(1-\dfrac{\left|{\mathbf{r}_{jk}}\right|}{\sigma}\right)^{2}&(\left|{\mathbf{r}_{jk}}\right|<\sigma)\cr 0&(\text{otherwise})\end{cases}$
 with very large barrier $V_{\text{max}}$, nearly equivalent to the hardcore interaction with diameter $\sigma$ (the particles are monodisperse). The inertia is completely ignored.
+Note that here we have adopted a system of monodisperse particles, which makes it simpler to specify the system.
+This is allowed as we are interested in the validity of the solid-based approach to only slightly glassy liquids.
 We prepare the system to be in a statistically steady and homogeneous state in equilibrium at the temperature $T$.
@@
 The system contains $N=4000$ particles, and $L$ is adjusted so as to give three values of the area fraction: $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$.
-To mimic the hardcore interaction, we chose $V_{\text{max}}=50$ for $\phi_{\text{area}}=0.50$, $V_{\text{max}}=500$ for $\phi_{\text{area}}=0.60$, and $V_{\text{max}}=5000$ for $\phi_{\text{area}}=0.70$ in units of ${k_{\text{B}}}T$; the time step for numerical integration, $\Delta{t}$, is chosen so as to satisfy $(\Delta{t}/\tau_{0})({k_{\text{B}}}T/V_{\text{max}})=0.1$ where $\tau_{0}=\sigma^{2}/D$.
+(For higher densities, polydisperse particles would be needed to avoid crystallization, which is out of the scope of the present work.) To mimic the hardcore interaction, we chose $V_{\text{max}}=50$ for $\phi_{\text{area}}=0.50$, $V_{\text{max}}=500$ for $\phi_{\text{area}}=0.60$, and $V_{\text{max}}=5000$ for $\phi_{\text{area}}=0.70$ in units of ${k_{\text{B}}}T$; the time step for numerical integration, $\Delta{t}$, is chosen so as to satisfy $(\Delta{t}/\tau_{0})({k_{\text{B}}}T/V_{\text{max}})=0.1$ where $\tau_{0}=\sigma^{2}/D$.
 In what follows, reference scales for nondimensionalization, such as $\sigma$ and $\ell_{0}$, will be shown explicitly as a rule.
@@
 Here we regard the elastic modeling as a convenient starting point for searching more sophisticated description of $C_{a}^{\Delta}$ with a wider range of applicability, involving solid-like elasticity at some scale.
-A recent experiment on confined liquid glyserol [39] may support relevance of such solid-like elasticity in liquids far away from glass transition.
-Given the numerical data of the DCs, we focus our attention to the ratio of the coefficients $S$ and $S/{\mu_{\text{r}}}$ in Eq.
+Recent experiments on confined liquid glyserol [43, 44] and polypropylene glycol [44] may support relevance of such solid-like elasticity in liquids far away from glass transition.
+Given the numerical data of the DCs, we focus our attention to the ratio of the coefficients $S$ and $S/{\mu_{\text{r}}}$ in Eqs.
 (51).
@@
 Using $f_{\pm}$ to express these correlations as
-$X_{\parallel}={\frac{1}{2}}\left({f_{+}}+{f_{-}}\right),\quad X_{\perp}={\frac{1}{2}}\left({f_{+}}+{f_{-}}\right)$ (56)
+$X_{\parallel}={\frac{1}{2}}\left({f_{+}}+{f_{-}}\right),\quad X_{\perp}={\frac{1}{2}}\left({f_{+}}-{f_{-}}\right)$ (56)
 in accordance with Eq. (41), we have
@@
 The analytical expressions of $f_{\pm}$ in Eqs.
-(61) and (62) allow us to extract information of elasticity from the numerical data of the particle system, on the assumption that the elastic model is quantitatively valid.
+(61) and (62) allow us to extract information of elasticity from simulational and experimental data of the particle system, on the assumption that the elastic model is quantitatively valid.
 They may provide useful alternatives to Eq.
@@
 It is easy to verify that the above expressions for DCs and SCs are consistently interrelated by Eqs.
-(45c).
+(45).
 In particular, Eq.
@@
 (76), which implies oscillatory divergence of the integrand for $k\to\infty$.
-Handling the divergence with the standard technique of convergence factor we find that $I_{a}^{(0)}(\xi_{*})$ vanishes for $\xi_{*}>0$ (the integral reduces to the delta function of $\xi_{*}$).
+Handling the divergence with the standard technique of convergence factor, we find that $I_{a}^{(0)}(\xi_{*})$ vanishes for $\xi_{*}>0$ (the integral reduces to the delta function of $\xi_{*}$).
 For $m=2$, using the convergence factor $e^{-\epsilon{k}}$ and taking the limit of $\epsilon\to+0$, we obtain
@@
 (12), from simulation data of a two-dimensional model colloidal liquid with $\phi_{\text{area}}=0.50$, $0.60$ and $0.70$.
+This is consistent with the experimental observation by Illing et al.
+[20].
 The time dependence of these correlations is not linear with regard to the time interval $t_{\Delta}$, as would be expected if the detected correlations were explained as accumulation of many events, but rather described in terms of a similarity variable $\vartheta$ in Eq.
@@
 The answer is negative: to answer this question, we have derived Eqs.
-(45c) as a relation between the two kinds of correlations, by treating the displacement field with the label variable formulation.
+(45) as a relation between the two kinds of correlations, by treating the displacement field with the label variable formulation.
 The computed values of DC and SC are then shown to be consistent with Eqs.
-(45c).
+(45).
 The relation in Eqs.
-(45c) takes the form of a linear mapping from the DC to the SC, which is expressed by means of two linear operators, $\hat{L}_{\pm}$, of Euler–Cauchy type.
+(45) takes the form of a linear mapping from the DC to the SC, which is expressed by means of two linear operators, $\hat{L}_{\pm}$, of Euler–Cauchy type.
 This mapping is non-invertible: we can discover $\chi_{\gamma}$ from the components of the DC, $f_{\pm}=X_{\parallel}\pm X_{\perp}$, but cannot recover $f_{\pm}$ uniquely from $\chi_{\gamma}$.
@@
 Thus we have evidenced that the DC is more informative than the SC. The extra information can be obtained from the numerical data by fitting Eq. (66) to the asymptotic logarithmic behavior of $f_{+}$.
-Interpretation of this information is our third question.
+Interpretation of this information, contained in $C_{+}$ and $\lambda$ obtainable through Eq.
+(66), is our third question.
 It is manifestation of the cage effect at spatiotemporal scales other than usually noticed.
@@
 (67), including the elastic modeling, can account for the logarithmic behavior of $f_{+}$ and the non-divergent behavior of $f_{-}$ at the shorter lengthscales.
+Thus the present work contributes to development of solid-based approach to liquid dynamics, allowing us to extract experimentally verifiable predictions about DC from theoretical modeling of $C_{a}^{\Delta}$, such as the elastic modeling.
 The elastic modeling, however, turns out to be too simple to reproduce the $\phi_{\text{area}}$-dependence of the coefficient $C_{+}$ of the logarithmic behavior in Eq.
-(66): while the numerical data suggests that $C_{+}$ is an increasing function of $\phi_{\text{area}}$, the elastic modeling makes the opposite prediction, as was noted toward the end of Subsec.
+(66): while the numerical data suggests (at least for $\phi_{\text{area}}\leq 0.7$) that $C_{+}$ is an increasing function of $\phi_{\text{area}}$, the elastic modeling makes the opposite prediction, as was noted toward the end of Subsec.
 VI.3.
@@
 Let us now remark on what could be done in the near future within the 2D setup.
-Since $C_{+}$ seems to carry information on some long-lived aspect of the cage effect at lengthscales greater than $\ell_{0}$, it will be desirable to develop more sophisticated modeling for the dynamics of $C_{a}^{\Delta}$ at such scales.
-Prediction of $C_{+}$ as a function of $\phi_{\text{area}}$ will serve as a touchstone of such modeling.
+First of all, we would like to invite experimentalists to measure the DC, verify Eq.
+(66) and thereby evaluate $C_{+}$.
+Since $C_{+}$ seems to carry information on some long-lived aspect of the cage effect at lengthscales greater than $\ell_{0}$, investigation of its dependence on $t_{\Delta}$ and $\phi_{\text{area}}$ by experiments (both real and simulated) will promote understanding of liquid dynamics at such scales.
+On the side of theoreticians, it will be desirable that they develop more sophisticated modeling for the dynamics of $C_{a}^{\Delta}$, which, with the aid of the formulae in Subsec.
+III.2, should allow prediction of $C_{+}$ as a function of $\phi_{\text{area}}$; its experimental verification will serve as a touchstone of such modeling.
 Viscoelastic modelings of mesoscopic deformations by combination of Maxwell-like phenomenology and the mode-coupling theory (MCT) would be a hopeful direction.
@@
 Details on this version of MCT will be reported elsewhere.
+Here we note, parenthetically, that MCT for tagged particles in Ref.
+[23] predicts DC with spatial oscillation unobserved in the experiment, and also that two completely different ideas for improvement on MCT for a tagged particle are discussed in Refs.
+[48, 30].
 As another possible direction to extend the present work, we may mention cases of higher $\phi_{\text{area}}$, which makes the liquid more glassy and the relaxation time long enough to allow comparison between the plateau regime, $\tau_{0}\ll t_{\Delta}\ll\tau_{\alpha}$, and the “post-$\tau_{\alpha}$” regime, $t_{\Delta}\gg\tau_{\alpha}$.
@@
 The first author (Ooshida) also thank Eni Kume and Shiladitya Sengupta for informative discussions on the occasion of recent international conferences held in the difficult time of the Covid19 pandemic.
-In addition, the assistance of Susumu Goto as a coauthor of Ref.
+Encouraging reactions to our arXiv preprint from Laurence Noirez and Alessio Zaccone are highly appreciated.
+Last but not least, the assistance of Susumu Goto as a coauthor of Ref.
 [24] is gratefully acknowledged, especially in preparation of Fig.